### v1

# Making Interpretable Discoveries from Unstructured Data: A High-Dimensional Multiple Hypothesis Testing Approach

Abstract
Social scientists are increasingly turning to unstructured datasets to unlock new empirical insights, e.g., estimating causal effects on text outcomes, measuring beliefs from open-ended survey responses. In such settings, unsupervised analysis is often of interest, in that the researcher does not want to pre-specify the objects of measurement or otherwise artificially delimit the space of measurable concepts; they are interested in “discovery.” This paper proposes a general and flexible framework for pursuing discovery from unstructured data in a statistically principled way. The framework leverages recent methods from the literature on machine learning interpretability to map unstructured data points to high-dimensional, sparse, and interpretable “dictionaries” of concepts; computes (test) statistics of these dictionary entries; and then performs selective inference on them using newly developed statistical procedures for high-dimensional exceedance control of the $k$-FWER under arbitrary dependence. The proposed framework has few researcher degrees of freedom, is fully replicable, and is cheap to implement—both in terms of financial cost and researcher time. Applications to recent descriptive and causal analyses of unstructured data in empirical economics are explored. An open source Jupyter notebook is provided for researchers to implement the framework in their own projects.
1 Introduction
Empowered by recent developments in machine learning and AI, researchers in the social sciences are increasingly leveraging sources of unstructured data—such as text, images, videos, and audio—in quantitative analyses. In economics, interest in unstructured data sources is especially widespread, ranging from using speech recordings from FOMC meetings to better understand monetary policy [Gorodnichenko et al., 2023]; to using videos of start-up pitches to study entrepreneurship and investment [Hu and Ma, 2025]; to using open-ended survey questions to probe the economic behavior and beliefs of individuals [Haaland et al., 2024]; to using mug shots to study judicial decisionmaking [Ludwig and Mullainathan, 2024]; to using visual art or written narratives to infer long-run living standards [Gorin et al., 2025, Lagakos et al., 2025]; to using interview transcripts to better understand the impacts of RCT treatments [Bergman et al., 2024, Krause et al., 2025].
The promise of these unstructured data, fundamentally, is that they offer new opportunities to measure social phenomena that were previously unmeasurable. This includes both measures that a researcher is capable of pre-specifying as well as those that a researcher is not, and instead would like to “discover.” When modern AI/ML methods are brought to bear on these goals, the former falls under the heading of “supervised learning,” whereby the researcher typically makes predictions of their ex-ante known quantity of interest using the unstructured data as features, and the latter falls under the heading of “unsupervised learning,” which is less amenable to a prediction-based framework, and demands that AI/ML methods uncover latent structure in the unstructured data with minimal researcher input. Though new statistical and econometric frameworks have recently become available for interpretable and statistically principled analyses of unstructured data in the supervised learning setting (e.g., Angelopoulos et al. [2023], Ludwig et al. [2024], Carlson and Dell [2025]), it is an open question as to how best perform interpretable and statistically principled analyses of unstructured data in unsupervised learning settings that emphasize discovery. The framework proposed in this paper provides one possible answer to this open question.
Specifically, this paper proposes a general framework for conducting interpretable and rigorous statistical inference on unstructured data in the following way, starting from an unstructured dataset $\{Z_{i}\}_{i=1}^{n}$:
  1. 1.
    Convert unstructured data to dictionary feature vectors: The researcher converts the unstructured data instances $Z_{i}\in\mathcal{Z}$ into high-dimensional, sparse, binary vectors $Y_{i}\in\{0,1\}^{p}$ using recently developed methods in dictionary learning for mechanistic interpretability of large language models (LLMs) and other deep neural networks (DNNs), i.e., $Y_{i}:=\mathtt{Dict}(Z_{i})$. Both theory and empirics corroborate that vector entries $Y_{ij}$ correspond to the presence of a monosemantic feature (i.e., a feature with a single, interpretable meaning) or “concept” $j$ in the given unstructured data instance $i$. There are $p\gg n$ dictionary features, e.g., we assume that $n$ may be on the order of $10^{2}$ or $10^{3}$, and $p$ is on the order of $10^{4}$ or $10^{5}$ (or larger). The breadth and depth of concepts spanned by these dictionary features is vast, as can be gleaned from the literature on dictionary learning for LLMs, and as will be observed concretely in the empirical applications of this paper. Intuitively, the dictionary should contain all possible (monosemantic) concepts that the LLM has learned for the purpose of next token prediction over a massive internet-scale corpus. However, due to sparsity imposed by design and suggested by machine learning theory, only a handful of these dictionary features will be present (“activate”) for any given unstructured data point. [^1] For more discussion of what “monosemanticity” means and its importance for LLM interpretability, see, e.g., the discussion in Bricken et al. [2023].
  2. 2.
    Compute test statistics from the dictionary feature vectors: The researcher computes from this newly created dictionary feature indicator dataset $\{Y_{i}\}_{i=1}^{n}$ a vector of $p$ test statistics, $T_{n}\in\mathbb{R}^{p}$. These test statistics are intended to test $p$ hypotheses, one for each dictionary feature, and may be flexibly chosen by the researcher. Especially important and simple choices explored in this paper’s empirical examples include hypotheses testing means and differences-in-means, feature by feature.
  3. 3.
    Perform high-dimensional selective inference using the test statistics: Finally, the researcher invokes a high-dimensional selective inference procedure newly developed in this paper to control the $k$-FWER of the selected set of rejected hypotheses (“discoveries”). This procedure is asymptotically valid for small $k$, requires only very mild assumptions on the data generating process (which are met in the applications of interest), and is well-powered in the empirical examples considered. This procedure may also be of independent interest in other high-dimensional selective inference settings.
The output of this framework is therefore a set of interpretable discoveries (rejections of dictionary feature derived tests) based on some set of unstructured data $\{Z_{i}\}_{i=1}^{n}$ with generalized familywise error rate guarantees, which may be relaxed or tightened as the researcher sees fit to permit more or fewer discoveries.
Two notable use cases for this framework explored in the empirical applications of this paper include:
  1. 1.
    Estimating average treatment effects on dictionary features in RCTs, or testing, for treatment indicator $D_{i}$ and fixed probability of treatment $\pi$,
    $H_{0,j}:E[Y_{ij}(1)-Y_{ij}(0)]=E[H_{i}Y_{ij}]=0,\quad H_{i}=\frac{D_{i}-\pi}{\pi(1-\pi)}.$
  2. 2.
    Estimating the probability of the presence of various concepts (measured by monosemantic dictionary features) in the population from which the unstructured dataset was sampled, or testing
    $H_{0,j}:E[Y_{ij}]=0$
    and then inverting these tests to form a “generalized” confidence set [Romano and Wolf, 2007].
This particular framework for discovery from unstructured data conveys several scientific benefits. The first is that of having few researcher degrees of freedom. In many settings featuring exploratory analysis of unstructured data, there is concern that researchers may cherry pick aspects of their data to measure or cherry pick choices of measurements in order to reverse engineer a specific conclusion. By design, however, the framework proposed by this paper has very little room for motivated data mining, as it requires no “human-in-the-loop” to direct what quantities are ultimately measured. Off-the-shelf, open-source dictionary learning models generate $p$ features spanning a huge number of concepts that are well-cataloged and publicly documented, making it unlikely the researcher needs to define additional concepts to measure ad hoc, and making a failure to report hypotheses for certain features conspicuous. Moreover, prior to analysis, a researcher could simply preregister a choice of dictionary learning model, such as particular sparse autoencoder (SAE) for a particular LLM, limiting the possibility that a researcher applies this framework with many different dictionary learning methods to select on favorable results. The proposed framework is also inexpensive both in terms of financial cost and researcher burden; both empirical examples in this paper were computed on Google Colab notebooks using a single A100 GPU, access to which only costs tens of dollars per month. [^2] This pricing description is valid as of October 2025. Further, this framework is easy to implement, requiring no human-in-the-loop to label unstructured data instances, no additional AI/ML model training, and no additional data sources; the framework is nearly automatic in execution, and takes only minutes to run using the above-mentioned hardware. This automaticity also lends the framework to fast and easy replication, another important scientific benefit.
Thus, the primary contributions of this paper are:
  1. 1.
    The development of a flexible, general-purpose framework for making interpretable as well as statistically and scientifically principled discoveries from unstructured data, leveraging state-of-the-art intepretability methods for LLMs and other DNNs.
  2. 2.
    The development of novel high-dimensional multiple hypothesis testing procedures for controlling $k$-FWER exceedance, facilitated by extensions of high-dimensional central limit theory.
The remainder of this paper is structured as follows: Section 2 discusses related literature; Section 3 describes the framework in detail; Section 4 applies the framework to two recent papers making use of unstructured data for discovery [Bursztyn et al., 2023, Stantcheva, 2024]; and Section 5 concludes.
2 Related Literature
The framework proposed in this paper is related to recent works from literatures spanning economics, statistics, and computer science.
Econometric methods for unstructured data.
Motivated by the capabilities of modern AI/ML methods for learning from unstructured economic data (see, e.g., Mullainathan and Spiess [2017], Gentzkow et al. [2019a], Dell [2025]), new econometric and statistical frameworks have been developed to facilitate principled statistical inference on low-dimensional features (predictions) learned from unstructured datasets under supervision, e.g., Angelopoulos et al. [2023], Ludwig et al. [2024], Carlson and Dell [2025], Rambachan et al. [2024]. This most recent wave of econometric literature emphasizes nonparametric frameworks compatible with black-box AI models coupled with debiasing methods, as opposed to model-based approaches, e.g., Gentzkow et al. [2019b], Battaglia et al. [2024]. Though the framework presented in this paper is concerned with principled inference on unstructured data without making parametric assumptions, it differs from this existing econometric literature in that its goal is primarily unsupervised discovery, as opposed to supervised detection, using AI/ML models.
Both the proposed framework and the recent econometric framework of Modarressi et al. [2025] support making causal inferences on text, with an emphasis on discovery. Modarressi et al. [2025] uses LLM reasoning as a tool for dimensionality reduction on the space of concepts present in text (i.e., as a low-dimensional hypothesis generation methodology) and incorporates human validation of LLM-selected hypotheses, with estimation implemented via sample splitting; in contrast, the framework proposed in this paper leverages recent advances in interpretability methods for LLMs in order to directly analyze the high-dimensional space of concepts present in text, and uses state-of-the-art autointerpretation methods to describe discovered concepts without further researcher intervention, and without requiring data splitting. As such, these frameworks offer different profiles of scientific benefits and researcher costs for the problem of causal inference on text.
Modarressi et al. [2025] also directly leverages the work of Ludwig et al. [2017] in their pipeline, which is similarly relevant to the proposed framework, in that both emphasize discovery from high-dimensional data, and can likewise both be applied to the problem of causal inference on unstructured data (as it is used in Modarressi et al. [2025]). However, the proposed framework tackles the high-dimensional discovery problem using a novel high-dimensional selective inference procedure that maintains a high degree of interpretability, whereas Ludwig et al. [2017] modifies the target of discovery to improve power, trading off interpretability to do so.
High-dimensional selective inference.
There is an extensive literature on multiple hypothesis testing (or “selective inference” more broadly) in statistics, biostatistics, and economics (see, e.g., Romano et al. [2010] for a review in econometrics). The methods in this literature span asymptotic and finite sample valid frameworks, and low- and high-dimensional settings. In particular, for the problem considered in this paper, the literature on asymptotically valid, high-dimensional selective inference is most relevant, as such methods permit making very few assumptions about the data (c.f., assuming independence or PRDS of the p-values in the case of Benjamini and Hochberg [1995]) and readily incorporate resampling methods that improve power (relative to guarding against worst-case dependence, c.f., Bonferroni corrections or the method of Benjamini and Yekutieli [2001]). These new procedures are built by extending recent work on high-dimensional central limit theory for order statistics [Chernozhuokov et al., 2022, Ding et al., 2025]. The proposed statistical procedures are the first stated of their kind that enable asymptotic high-dimensional exceedance control of the $k$-FWER for test statistics that may be arbitrarily correlated with one another, complementing existing methods for asymptotic high-dimensional control of the FWER [Belloni et al., 2018] and the FDR under weak dependence [Liu and Shao, 2014, Belloni et al., 2018].
Hypothesis generation.
There is a growing literature in economics and computer science on “hypothesis generation” at the pre-scientific stage of empirical inquiry [Ludwig and Mullainathan, 2024]. From the perspective of this literature, the proposed framework may be viewed as leveraging dictionary learning as a high-dimensional hypothesis generator, on top of which discovery is performed at the “scientific” stage of inquiry using high-dimensional multiple hypothesis testing procedures. As such, the framework in this paper contributes to the literature on hypothesis generation by providing a statistically principled pathway from generation to evaluation—one which involves little to no researcher discretion about what generated hypotheses should be tested. The recent framework of Movva et al. [2025] uses dictionary learned features (specifically, sparse autoencoder features) explicitly for the purposes of hypothesis generation, further motivating the use of dictionary learning methods as tools for hypothesis generation in the social sciences, both in the present framework and beyond.
Interpretability for LLMs and other DNNs.
There is a large literature in computer science on machine learning interpretability (see, e.g., Doshi-Velez and Kim [2017] for a foundational agenda). The machine learning interpretability methods leveraged by this framework originate from a nascent though highly active literature known as “mechanistic interpretability,” which seeks to develop methods for interpreting LLM behavior via quantitative analyses of model internals (e.g., activations, weights). In particular, the dictionary learning methods for DNN interpretability implemented in this framework were developed in Bricken et al. [2023], Templeton et al. [2024], i.e., sparse autoencoders (SAEs). SAEs are autoencoders attached to the residual streams (or other model internals) of a LLM, with hidden layers being many orders larger in dimension than the residual stream, and which are trained with sparsity-inducing penalization on reconstruction loss. Under the “linear representation hypothesis” and “superposition hypothesis” (see, e.g., Bricken et al. [2023] for more discussion), these sparse autoencoders are thought to act as an overcomplete basis of the space of concepts in text leveraged by a LLM to make next token predictions, encouraging learned features to be monosemantic. There has been much debate about the success of SAEs (e.g., Leask et al. [2025]), though most criticisms still support the notion that SAEs and other dictionary learning methods have a comparative advantage in discovering as opposed to detecting concepts of interest in text [Peng et al., 2025]. Other dictionary learning methods that have been proposed as competitors to SAEs (e.g., transcoders [Paulo et al., 2025]) are similarly compatible with the framework proposed in this paper.
SAEs and other dictionary learning methods generate features, but not natural language feature descriptions. Based on the popularity of these methods, a related literature on “autointerpretability” methods has become active, which seeks to coherently use LLMs to describe the features discovered by dictionary learning methods at scale. Recent important papers in this literature include: Bills et al. [2023], Shaham et al. [2024], Paulo et al. [2024], Rajamanoharan et al. [2024]. The framework proposed in this paper primarily uses insights from Paulo et al. [2024] to create “local interpretations” that are most relevant to the data generating process being considered, and otherwise leverages autointerpetations from Lieberum et al. [2024].
Though the empirical examples considered in this paper handle one of the most popular unstructured date types—text—for which dictionary learning-based interpretability methods are the most well-developed, dictionary learning techniques have been successfully applied to many other modalities, including audio and images [Abdulaal et al., 2024, Bhalla et al., 2024, Fry, 2024, Daujotas, 2024, Pluth et al., 2025].
3 Framework
3.1 Setup
Let $n]:=\{1,\dots,n\}$, and let $x\mapsto\log x$ be the natural logarithm. The researcher has access to a dataset of size $n$, $\{(W_{i},Z_{i})\}_{i=1}^{n}$, which is sampled i.i.d. from some (super-)population of interest $P$. The $Z_{i}\in\mathcal{Z}$ are unstructured data instances and the $W_{i}\in\mathcal{W}$ are any other observed covariates of interest (e.g., for one empirical example considered in the following section, $W_{i}=D_{i}\in\mathcal{W}=\{0,1\}$, a binary treatment status indicator). The space $\mathcal{Z}$ is typically high-dimensional and semantically poor, e.g., if each $Z_{i}$ was a $244\times 244$ pixel image, $\mathcal{Z}$ might then be the space of all $3\times 244\times 244$ arrays of RGB values, or if $Z_{i}$ was text that was truncated up to some maximum length, $\mathcal{Z}$ might be the space of all binary matrices of a certain dimensionality, which are concatenated one-hot encodings for each word, term, or token in the text with respect to a specific pre-defined vocabulary.
The researcher has access to the function $\texttt{Dict}:\mathcal{Z}\to\{0,1\}^{p}$, which maps an unstructured data observation to a a sparse binary vector. This function is computed by passing an unstructured data point to a DNN equipped with a dictionary learning model (e.g., a LLM equipped with pretrained SAE when $Z_{i}$ are texts), recording the feature activations from the dictionary model, and then aggregating and binarizing these dictionary feature activations to form a single indicator for each feature in the dictionary $j\in[p]$. There is a large possible space of Dict functions the researcher could implement, though an especially straightforward choice for the use case of text data is implementing a Dict function based on the SAE activations at a single layer of the DNN, for which the $j$-th entry of the output is equal to 1 if the $j$-th dictionary feature activated on any token of the input text, and 0 otherwise:
[…]
We denote $Y_{i}:=\mathtt{Dict}(Z_{i})\in\{0,1\}^{p}$ as the output of the dictionary transformation of the unstructured data. As such, each $Y_{ij}\in\{0,1\}$ has the natural interpretation that a particular dictionary feature $j$ activated for the unstructured data instance $i$, and, e.g., $\bar{Y}_{\bullet,j}:=n^{-1}\sum_{i=1}^{n}Y_{ij}$ corresponds to an estimate of the probability of the presence of a feature $j$ in the population $P$ of interest. [^3] Dictionary feature activations are typically positively valued scalars, where magnitude is thought to correspond to some notion of intensity of activation; other works in mechanistic interpretability consider using these raw activation values directly, perhaps with max or average pooling across tokens. However, presently, magnitude of activation is not a well-understood or highly interpretable quantitative property of dictionary learning methods, and as such the default Dict function advocated for in the proposed framework does not incorporate this information. That said, researchers interested in extending this framework may find value in alternative Dict functions that leverage activation intensity.
The crux of the proposed framework is that inference on functionals of $Y_{ij}$ is desirable because:
  1. 1.
    Each $Y_{ij}$ is monosemantic, and therefore functionals of $Y_{ij}$ for a given $j$ are interpretable, e.g., $E_{P}[Y_{ij}]$ is the probability of monosemantic feature $j$ appearing in an unstructured data observation in population $P$.
  2. 2.
    Collectively, the $p$ features in $Y_{ij}$ for any given $i$ are exhaustive of some vast space of concepts of interest to the researcher; intuitively, they are the set of all monosemantic features that a LLM needed to learn to perform well on next token prediction for a massive, internet-scale corpus of text (or other similarly massive unstructured data source). This means the researcher need not pre-specify any particular concepts of interest, reducing researcher degrees of freedom without limiting discovery.
However, inference on functionals of $Y_{ij}$ for each $j$ is also challenging, because:
  1. 1.
    The $p$ is large, so inference on all $j\in[p]$ functionals is inherently a multiple hypothesis testing problem.
  2. 2.
    Not only is $p$ large, but for most social science applications of interest $p\gg n$, so inference on $Y_{ij}$ for each $j$ is a high-dimensional multiple hypothesis testing problem.
  3. 3.
    Each estimator, test statistic, or p-value formed from the $Y_{ij}$ for each $j\in[p]$ for the purposes of inference is plausibly statistically dependent on every other in a complicated way, ruling out multiple hypothesis testing approaches that assume independence or specific forms of dependence (e.g., PRDS).
  4. 4.
    Each $Y_{i}$ is sparse, and the intent of the analysis is discovery, so the desired form of selective (familywise) error control cannot be too conservative.
What is needed, then, for principled inference on functionals of $Y_{ij}$ is a high-dimensional selective inference procedure with control over a generalized error rate. To be as well-powered as possible, we want to focus on testing procedures that employ resampling methods for estimating the true covariance across test statistics, such that conservative protection against worst-case dependence is not required. To keep things as general as possible, we also want to allow for only asymptotically valid test statistics and p-values, as is common to much of econometric analysis. In the following sections, theory and corresponding statistical procedures are developed to achieve exactly these aims, and are stated as generally as possible to accommodate settings of interest even beyond analysis of unstructured data.
To pursue inference, consider defining $p\gg n$ “one-sided” hypotheses $\{H_{0,j}\}_{j\in[p]}$, where
$H_{0,j}:\theta_{j}(P):=E_{P}[X_{ij}]\leq 0,\quad X_{ij}=h(W_{i},Y_{ij}),$
for some measurable function $h$. Two choices of $h$ considered in the empirical examples in this paper include $h(W_{i},Y_{ij})=Y_{ij}$ (testing probabilities of dictionary feature activation) and $h(W_{i},Y_{ij})=\frac{D_{i}-\pi}{\pi(1-\pi)}Y_{ij}$ (testing differences in probabilities of activation across groups under known, fixed group assignment), though many others are possible, lending to the flexibility of the framework. As shorthand, we will denote $X:=\{X_{i}\}_{i=1}^{n}$. Consider also defining a set of “two-sided” hypotheses of interest $\{\tilde{H}_{0,j}\}_{j\in[p]}$ as
$\tilde{H}_{0,j}:\theta_{j}(P):=E_{P}[X_{ij}]=0.$
We may form test statistics for such hypotheses as
$T_{n,j}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{ij},\quad T_{n}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{i}.$
For the purposes of developing high-dimensional central limit theory later on, further define
$S_{n,j}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{ij}-E[X_{ij}]),\quad S_{n}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-E[X_{i}]).$
We will denote $\Sigma:=E\left[S_{n}S_{n}^{\mathtt{T}}\right]=n^{-1}\sum_{i=1}^{n}E[(X_{i}-E[X_{i}])(X_{i}-E[X_{i}])^{\mathtt{T}}]$, which simplifies to $\Sigma=E[(X_{i}-E[X_{i}])(X_{i}-E[X_{i}])^{\mathtt{T}}]$ under i.i.d. data. Further, for any subset of indices $K\subseteq[p]$, we will denote $S_{n,K}:=(S_{n,j}:j\in K)^{\mathtt{T}}$.
3.2 High-Dimensional $k$-FWER Exceedance Control
We construct methods that provide high-dimensional $k$-FWER exceedance control asymptotically, or control of the probability of making $k$ or more rejections of true nulls in the large sample limit when $p$ is growing (much) faster than $n$ in the relevant asymptotic thought experiment. Naturally, $k$-FWER control for $k=1$ is control of the FWER, and larger choices of $k$ permit more discoveries.
To proceed, we will adapt strategies from Romano and Wolf [2007] for achieving large sample $k$-FWER control to the high-dimensional setting. The testing procedures of Romano and Wolf [2007] rely on the asymptotic validity of bootstrap approximations of the distribution of the $k$-th largest coordinate (or “$k$-max”) of $S_{n}$. Thus, a bootstrap procedure for the $k$-max statistic that is asymptotically valid in high-dimensions would allow one to run the algorithms of Romano and Wolf [2007] in the high-dimensional setting with their stated control under minimal modifications.
High-dimensional bootstraps must be supported by appropriate high-dimensional central limit theory. Fortunately, the recent work of Ding et al. [2025] provides high-dimensional central limit theory for the $k$-th largest coordinate of a scaled sum of centered independent random vectors so long as $k$ is very small, i.e., $k$ is fixed in the asymptotic thought experiment. We will build on these “small kk” results. Assume $n\geq 3$ and $p\geq 3$. Let $b_{1}>0$ and $b_{2}>0$ be some constants such that $b_{1}\leq b_{2}$, and let $B_{n}>1$ be a sequence of constants, where it is possible that $B_{n}$ diverges. Then we make the following assumptions.
Assumption M.
For all $i\in[n],j\in[p]$, assume: (i) $E\left[\exp\left(\left|X_{ij}\right|/B_{n}\right)\right]\leq 2$; (ii) $b_{1}^{2}\leq\frac{1}{n}\sum_{i=1}^{n}E[X_{ij}^{2}]$; and (iii) $\frac{1}{n}\sum_{i=1}^{n}E[X_{ij}^{4}]\leq B_{n}^{2}b_{2}^{2}$.
These are mild conditions on the tails and moments of the data $X_{ij}$, which are stated in a way that accommodates data that are independent but not identically distributed (i.n.i.d.). Part (i) of Assumption M simply requires that $X_{ij}$ be sub-exponential, or, equivalently, have an Orlicz norm in $\Psi_{1}$ bounded by $B_{n}<\infty$. The $B_{n}$ is indexed by $n$ to accommodate growing tail thickness within the sub-exponential regime as $p=p_{n}$ grows in the asymptotic thought experiment with i.n.i.d. data. Part (ii) of Assumption M insists that the second moments of the data be bounded away from zero, appropriately stated for i.n.i.d. data. Part (iii) insists on bounded fourth moments in a similar fashion. As such, for i.i.d. data, the case considered in this paper, we may simply require for some fixed $B<\infty$ and fixed $b_{1},b_{2}$ with $b_{1}\leq b_{2}$: (i) $E\left[\exp\left(\left|X_{ij}\right|/B\right)\right]\leq 2$; (ii) $b_{1}^{2}\leq E[X_{ij}^{2}]$; and (iii) $E[X_{ij}^{4}]\leq B^{2}b_{2}^{2}$. Note that, in many use cases of interest for dictionary learned features, the $X_{ij}$ are both bounded and studentized, aiding the plausibility of these assumptions.
In practice, some fraction of dictionary features may be “dead” for any given dataset, meaning they are degenerate at zero [Bricken et al., 2023]. Of course, these features are not of interest for the purposes of discovery. This phenomenon is tolerated by the framework so long as the count of degenerate features is smaller than any relevant pkp-k, in which case, in any finite sample, the $k$-th largest coordinate of all relevant statistics is invariant to whether or not the degenerate coordinates are dropped. As such, conceptually, one should think of $j$ as ranging over the $p$ non-degenerate features. In practice, to perform studentization, it is recommended that these degenerate coordinates are explicitly filtered out; studentization is discussed further in Section 3.3.
Assumption R.
Assume that $B_{n}^{2}\log^{5}(pn)=o(n)$.
This is the key rate condition needed for (sup-norm) Gaussian or bootstrap approximation error to go to zero asymptotically under the high-dimensional CLTs discussed in both Ding et al. [2025] and Chernozhuokov et al. [2022]. Rewritten, it says that
$\frac{B_{n}^{2}\log^{5}(pn)}{n}=o(1)$
which permits $p$ growing very fast with $n$ in the asymptotic thought experiment. In fact, $p$ may be growing nearly exponentially in $n$, e.g., $p=e^{n^{1/6}}$ for fixed $B_{n}$. That this rate condition permits $p\gg n$ is an important positive result in high-dimensional central limit theory [Chernozhukov et al., 2017, 2023].
We will focus on the Gaussian multiplier bootstrap as our high-dimensional bootstrap method. We define the Gaussian multiplier bootstrap quantity
$S_{n}^{B}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\xi_{i}\left(X_{i}-\bar{X}_{n}\right),$
where $\xi_{i}\overset{iid}{\sim}N(0,1)$ and $\bar{X}_{n}:=n^{-1}\sum_{i=1}^{n}X_{i}$. We also introduce $x\mapsto x_{[k]}$ as notation for the function that selects the $k$-th largest coordinate of a vector $x$, i.e., $x_{[k]}=k\text{-}\!\max(x)$ in the notation of Romano and Wolf [2007]. For the purposes of bootstrapping, we will denote the data-conditional probability measure $P^{B}(\cdot):=P(\cdot\mid X)$.
With these assumptions and notations in place, we now introduce statistical procedures for controlling the $k$-FWER for small $k$, based on the step-wise algorithms of Romano and Wolf [2007].
{restatable}
[High-dimensional $k$-FWER exceedance control for small $k$, one-sided]theoremSmallk Consider the method of Algorithm 2.1 or 2.2 in Romano and Wolf [2007] with test statistics $T_{n}$ of hypotheses $\{H_{0,j}\}_{j\in[p]}$ and critical values $\hat{c}_{n,K}(1-\alpha,k)$ given by the $1-\alpha$ quantile of $S_{n,K,[k]}^{B}$ under $P^{B}$. Assume that $k$ is fixed (i.e., not growing with $n,p$). Then under Assumptions M and R:
  1. (i)
    $\limsup_{n,p\to\infty}k\text{-FWER}_{P}\leq\alpha$.
  2. (ii)
    If $H_{0,j}$ is false and $\theta_{j}(P)\gg B_{n}\sqrt{\log p/n}$, then the probability that the step-down method rejects $H_{0,j}$ tends to 1.
{restatable}
[High-dimensional $k$-FWER exceedance control for small $k$, two-sided]theoremSmallktwo Consider the method of Algorithm 2.1 or 2.2 in Romano and Wolf [2007] with test statistics $|T_{n}|$ of hypotheses $\{\tilde{H}_{0,j}\}_{j\in[p]}$ and critical values $\hat{c}_{n,K}(1-\alpha,k)$ given by the $1-\alpha$ quantile of $|S_{n,K}^{B}|_{[k]}$ under $P^{B}$. Assume that $k$ is fixed (i.e., not growing with $n,p$). Then under Assumptions M and R:
  1. (i)
    $\limsup_{n,p\to\infty}k\text{-FWER}_{P}\leq\alpha$.
  2. (ii)
    If $\tilde{H}_{0,j}$ is false and $|\theta_{j}(P)|\gg B_{n}\sqrt{\log p/n}$, then the probability that the step-down method rejects $\tilde{H}_{0,j}$ tends to 1.
Importantly, the procedures of Theorems R and R are only valid for small $k$. To see this explicitly, note that the sup-norm bootstrap or Gaussian approximation error in the high-dimensional CLTs for the $k$-th largest coordinate introduced in Ding et al. [2025] only goes to zero if
$\frac{k^{8}B_{n}^{2}\log^{5}(pn)}{n}=o(1),$
meaning that $k$ must be fixed or must grow incredibly slowly: for all practical purposes, $k$ needs to be quite small. As such, the theory of Ding et al. [2025] is unsuitable for making progress on FDP exceedance control, which requires a $k$ that may grow linearly with $p$. An important goal of future research would be to establish valid high-dimensional FDP exceedance control using appropriate high-dimensional central limit theory.
3.3 High-Dimensional CLT for $k$-max of Approximate Means
In practice, hypothesis testing with studentized statistics will be important for powering discoveries beyond just frequently occurring dictionary features (which are, e.g., high variance features for Bernoulli random variables when $X_{ij}=Y_{ij}\in\{0,1\}$). To facilitate studentization, we will need a high-dimensional CLT that permits statistics with small estimation errors. We therefore generalize the high-dimensional CLTs of Ding et al. [2025] to handle “approximate means” in the parlance of Belloni et al. [2018], stated in the following lemmas.
{restatable}
[High-dimensional CLT for the small $k$-max coordinate of approximate means]lemmaHdapprox Let $\hat{S}_{n}:=S_{n}+R_{n}$, and assume that $\|R_{n}\|_{\infty}=o_{P}(1/\sqrt{\log p})$. Further assume that $k$ is fixed (i.e., does not grow with $n,p$). If Assumptions M and R hold, then as $n,p\to\infty$
$\sup_{t\in\mathbb{R}}\left|P\left(\hat{S}_{n,[k]}\leq t\right)-P\left(N(0,\Sigma)_{[k]}\leq t\right)\right|\to 0.$
{restatable}
[High-dimensional bootstrap for the small $k$-max coordinate of approximate means]lemmaHdapproxboot Let $\hat{S}_{n}^{B}:=S_{n}^{B}+R_{n}$, and assume that $\|R_{n}\|_{\infty}=o_{P}(1/\sqrt{\log p})$. Further assume that $k$ is fixed (i.e., does not grow with $n,p$). If Assumptions M and R hold, then as $n,p\to\infty$
$\sup_{t\in\mathbb{R}}\left|P^{B}\left(\hat{S}_{n,[k]}^{B}\leq t\right)-P\left(N(0,\Sigma)_{[k]}\leq t\right)\right|\xrightarrow[]{P}0.$
Using these lemmas, we may now state two corollaries, which allow for studentization. First, define $\Lambda:=\text{diag}(\Sigma)$, as well as $\Sigma_{0}:=\Lambda^{-1/2}\Sigma\Lambda^{-1/2}$. Let $\widehat{\Sigma}_{jj}:=n^{-1}\sum_{i=1}^{n}\left(X_{ij}-\bar{X}_{n,j}\right)^{2}$ and let $\widehat{\Lambda}:=\operatorname{diag}\left\{\widehat{\Sigma}_{11},\ldots,\widehat{\Sigma}_{pp}\right\}$ be an estimator of the asymptotic variances $\Lambda$.
{restatable}
[High-dimensional CLT for the small $k$-max studentized coordinate]corollaryCorhd If Assumptions M and R hold, and $B_{n}=O(1)$ and $k$ is fixed (i.e., does not grow with $n,p$), then, by application of Lemma 3.3, as $n,p\to\infty$
$\sup_{t\in\mathbb{R}}\left|P\left(\left(\hat{\Lambda}^{-1/2}S_{n}\right)_{[k]}\leq t\right)-P\left(N(0,\Sigma_{0})_{[k]}\leq t\right)\right|\to 0.$
{restatable}
[High-dimensional bootstrap for the small $k$-max studentized coordinate]corollaryCorhdboot If Assumptions M and R hold, and $B_{n}=O(1)$ and $k$ is fixed (i.e., does not grow with $n,p$), then, by application of Lemma 3.3, as $n,p\to\infty$
$\sup_{t\in\mathbb{R}}\left|P^{B}\left(\left(\hat{\Lambda}^{-1/2}S_{n}^{B}\right)_{[k]}\leq t\right)-P\left(N(0,\Sigma_{0})_{[k]}\leq t\right)\right|\xrightarrow[]{P}0.$
With these corollaries, it is immediate that the bootstrap quantiles of studentized statistics well approximate the studentized version of $S_{n}$. Also note that, in the setting where the data is not just i.n.i.d. but i.i.d., as is true for the use case of interest based on dictionary features $X_{i}$, $B_{n}=O(1)$ is assured.
4 Empirical Applications
As an illustration of the framework proposed in this paper, we now reanalyze two recent works in empirical economics that pursue discovery from unstructured data, and show how new, principled, and interpretable discoveries may be made using the same exact data at low cost.
4.1 Bursztyn et al. [2023]
Bursztyn et al. [2023] study how the provision of “social cover” affects willingness to publicly dissent to socially stigmatized causes, and the perception of this dissent. As a key application of their theory, Bursztyn et al. [2023] run an information treatment experiment (Experiment 2) in which participants are told they have been matched with another (fictional) respondent that chose to join a campaign to defund the police (a plausible expression of dissent in liberal American politics at the time the experiment was conducted), and then show the participant a tweet that the matched respondent is said to have agreed to publicly post. This tweet has been randomized to either include social cover or no social cover, namely whether or not the tweet indicated that the matched respondent joined the campaign after reading an evidence-based article in support of it (the no social cover condition) or before (the social cover condition). [^4] See Online Appendix Figure B.3 of Bursztyn et al. [2023] for a schematic of the experimental design.
An important outcome collected by this experiment is the participant’s open-ended text response to the question “Why do you think your matched respondent chose to join the campaign to oppose defunding the police?” This open-ended text response is meant to capture the causal effect of social cover on the perception of dissent. As such, this is a setting where discovery is of interest; ideally, to form a holistic understanding of how social cover affects perception of dissent, we do not want to pre-specify what aspects of perception we are interested in, and want to discover any interpretable systematic differences that exist in the text responses across treatment and control groups. In order to make progress on this analysis of unstructured data, Bursztyn et al. [2023] compute a Pearson’s $\chi^{2}$ statistic for all phrases of up to three words per Gentzkow and Shapiro [2010], which they use as an index to rank the phrases that are most differentially expressed in each condition’s open-ended responses. [^5] Specifically, it is a Pearson’s $\chi^{2}$ statistic for a null hypothesis that the propensity to use a given phrase is equal across conditions, per Gentzkow and Shapiro [2010]. We interpret these statistics as simply forming an index, however, because the results of these hypothesis tests are not reported, and no multiple hypothesis testing corrections are implemented. The interpretability of the results from this analysis, is, naturally, hindered by the coarseness of the featurization of the text as three word phrases, as well as the fact that there are no obvious estimands or inferential guarantees. The only qualitative conclusion gleaned by Bursztyn et al. [2023] from this quantitative exercise is: “we find that respondents in the Cover condition are more likely to use phrases related to the article or the associated evidence—for example, ‘article,’ ‘read,’ ‘convincing,’ or ‘increase in crime.’ ” [^6] The Online Appendix Table B.11 of Bursztyn et al. [2023] contains the top ten characteristic phrases in each condition based on the $\chi^{2}$ index.
Can we make more (and more interpretable) discoveries with the framework of this paper? To investigate, we implement a Dict function based on the Gemma Scope autoencoders for Google’s Gemma 2 2B [Lieberum et al., 2024], and specifically the SAE trained on the residual stream of layer 12.777It is recommended to work with SAEs trained on the middle layer of a LLM to capture coarser concepts and features, which is plausibly more suitable for analyzing broad themes in text. We pull the Gemma Scope project’s autointerpretation descriptions of these features from Neuronpedia, [^8] See https://www.neuronpedia.org/gemma-scope for more. though we also implement a custom, “local” autointerpretation pipeline based on the best practices of Paulo et al. [2024], discussed further in Appendix Section 6.2. These are mainstream choices for a SAE and LLM pairing in machine learning interpretability research.
This Dict function yields a dataset $\{X_{ij}\}_{i\in[n],j\in[p]}$, where $n=1033$ and p=12,005p=12,\!005 (non-degenerate features). The hypotheses of interest are, for all $j\in[p]$,
$H_{0,j}:E[Y_{ij}(1)-Y_{ij}(0)]=E[H_{i}Y_{ij}]:=E[X_{ij}]=0,\quad H_{i}=\frac{D_{i}-\pi}{\pi(1-\pi)}.$
for $\pi=0.5$, and where $\{D_{i}=1\}$ indicates the social cover condition. As a baseline, we will implement a simple one-step version of the statistical procedure in Theorem R, which will attain a lower bound on the power achieved by the step-wise procedure.
We may first consider whether or not, as anticipated, simply controlling the FWER would be too conservative for the purposes of discovery in this setting. To explore this, we set $k=1$ and $\alpha=0.05$ in the procedure based on Theorem R (equivalently, we implement the high-dimensional FWER controlling procedure of Belloni et al. [2018]), and we obtain only two significant discoveries: a 13.2 percentage point causal effect on the presence of dictionary feature 3518 (DF-3518) and a 13.4 percentage point causal effect on the presence of dictionary feature 3426 (DF-3426). The Gemma Scope autointerprations of these features paint a picture that corroborates the original findings of Bursztyn et al. [2023]: DF-3426 is described as activating on “references to articles” and DF-3426 on “references to news articles and reports.” However, we only discover two features out of over 12,00012,\!000, and not much new insight relative to the original analysis—even if the interpretation of the features discovered was transparent and automatic, by construction.
In Table 1, we make discoveries based on the procedure of Theorem R now setting $k=5$, still with $\alpha=0.05$, and this time using local autointerpretation descriptions per the method Appendix Section 6.2.
Table 1: Discoveries for Experiment 2 of Bursztyn et al. (2023), 5-FWER Control at $\alpha=5\%$
Feature $\widehat{\text{ATE}}$ Description (“Feature activates highly on…”) t-stat
306 0.099 source that claims police defunding increases crime 4.103
1392 0.118 partial or incomplete reading of an article 3.910
1622 0.072 mentions of a prestigious author or institution (e.g., “Princeton professor”) 4.045
1992 0.095 mentions of an article 4.286
3298 0.126 article about defunding the police and its impact on violent crime 4.223
3426 0.134 mentions of an “article” or article reference 4.803
3518 0.132 articles linking police defunding to increased violent crime 5.020
4320 0.081 article or report content 3.999
8486 0.085 defunding‑the‑police political discourse 3.918
8902 0.079 reading an article that influences a person’s beliefs or viewpoints 4.198
12287 -0.130 modal verbs expressing uncertainty or possibility 3.890
Note: Descriptions generated using the method described in Appendix Section 6.2, using GPT-OSS 20B [OpenAI, 2025]. For bootstrapping, 10,00010,\!000 iterations are performed.
As can be seen, even with control of $k$-FWER at $k=5$, we obtain 11 rejections of the null, or 11 discoveries. By guarantee of Theorem R, the probability that 5 or more of these 11 discoveries is false is less than 5%. The previous features discovered with the FWER controlling procedure appear, as well as others that represent concepts both absent from the FWER controlled analysis and the original analysis in Bursztyn et al. [2023]. Though there are features that appear to activate on redundant concepts, in totality these discoveries contribute to entirely new characterizations of the causal effect of social cover on dissent. For example, DF-1622 responds to “mentions of a prestigious author or institution” and increases by 7 percentage points in the treatment group, indicating that, as Bursztyn et al. [2023] posit, the social cover mechanism plausibly relies on the perceived credibility of the article cited. DF-1392 increases by 12 percentage points in the treatment group, and responds to discussion of “partial or incomplete reading of an article,” suggesting that that social cover in this experiment may have a negatively valenced effect related to perceptions of ignorance. DF-12287 responds to “modal verbs expressing uncertainty or possibility,” and its probability of activation is decreased by 13 percentage points in the treatment group. This supports the idea that the causal effect of not having social cover is, to first order, generating uncertainty and speculation about dissent, as opposed to an alternative where participants coalesce around a single narrative in the absence of cover. This insight is also congruous with the observation that few large negatively signed causal effects are discovered.
Importantly, all of this analysis was performed automatically, in a handful of minutes on a single A100 GPU in Google Colab. Though this analysis was not preregistered, it is possible to cheaply and automatically replicate these results, and assess their sensitivity to different autointerpretation strategies, dictionary learning methods, and LLMs. Notably, the space of possible concepts was not pruned in any way prior to this analysis, nor were they weighted in some way towards topics related to articles; the analysis was automatic and without human intervention, with no room for motivated data mining.
4.2 Stantcheva [2024]
Stantcheva [2024] conducts surveys on representative samples of the United States population investigating attitudes towards inflation. These surveys include open-ended text responses to a variety of questions, with the goal of discovering attitudes and opinions that the researcher should not (for concern about priming respondents) or could not (for lack of imagination) pre-specify. [^9] See, e.g., Haaland et al. [2024] for more discussion of the benefits of open-ended survey questions for understanding economic behavior.
An important open-ended prompt that Stantcheva [2024] solicit an answer to is “High inflation is caused by…” In the original analysis, Stantcheva [2024] code open-ended responses with a keyword-based method described in Ferrario and Stantcheva [2022], in which the researcher creates a list of topics and associated keywords in a discretionary way, ranging from “manual to semi-supervised or unsupervised” means. The result of applying this analysis to the open-ended responses to the above prompt yields Figure 3 in Stantcheva [2024], which finds, among other things, that mentions of “Biden and the administration,” “Greed,” “Monetary policy,” “Fiscal policy,” “War and foreign policy,” “Demand vs supply,” and “Supply-side mechanisms (other than input prices)” appear in more than 5% of all responses.
Do we make the same discoveries automatically if we apply the proposed framework instead? Do we make more and new interpretable discoveries? To find out, we use the same models as in Section 4.1, but reduce the space of features by half, filtering out all features that, in the corpus on which the Gemma Scope SAEs were trained, had greater than median empirical token activation frequency. This choice of dimensionality reduction is meant to screen out features that activate in many texts across domains, e.g., features related to grammatical aspects of text, which was handled implicitly in the previous causal analysis by virtue of differencing. (Ideally, such a dimensionality reduction choice would be preregistered, to prevent cherry-picking by filtering.) This yields a dataset $\{X_{ij}\}$ with $n=503$ and $p=3915$ non-degenerate features—still a very high-dimensional inference problem.
We test the $p=3915$ hypotheses $H_{0,j}:E[X_{ij}]=E[Y_{ij}]=0$, i.e., we test the probability of a given feature activating in the population of open-ended responses to the above prompt in the United States. Specifically, by again using 5-FWER control with $\alpha=0.05$ based on the procedure of Theorem R, we make 818 interpretable discoveries. In order to better present these discoveries, using test inversion we produce simultaneous “generalized” confidence sets [Romano and Wolf, 2007], which asymptotically guarantee that the probability that $k$ or more estimands do not lie in the set is less than $\alpha$.
In Table 2, we report the simultaneous confidence intervals for all discoveries with a lower confidence bound above 20%. Notably, we discover, automatically, with no manual pruning or interpretation required, many sensible commonplace themes related to the topic at hand: DF-9804 responds to economic terms, DF-14304 responds to economic terminology, and DF-13447 responds to mentions of inflation. However, we also recover many of the topics from the original analysis in Stantcheva [2024] without any human discretion or intervention: DF-4192 responds to discussion of monetary policy, DF-5719 responds to political commentary, and DF-14747 responds to discussion of supply-side issues. Undoubtedly, many other topics are recovered as we look deeper down the list of discoveries, beyond the top 11 out of 818. However, even in this top 11, we learn that there is a great deal of pure uncertainty expressed in these open-ended responses based on the activation of DF-4794; that macroeconomic indicators get frequently discussed by virtue of DF-11036; and that money supply and government spending and budgeting are often mentioned, per DF-104 and DF-8316, respectively.
Table 2: Largest Discoveries for “High inflation is caused by…” in Stantcheva (2024)
Feature Description (“Feature activates highly on…”) Sim. CI, Lower Sim. CI, Upper
104 excess money supply 0.208 0.352
4192 monetary policy and central‑bank terminology 0.217 0.363
4794 expressions of uncertainty or hedging 0.210 0.355
5719 political commentary about government or the economy 0.843 0.942
8316 government budget or spending 0.279 0.433
9804 economic terms related to the economy 0.268 0.420
11036 macroeconomic indicators 0.212 0.357
13447 inflation or rising prices 0.393 0.553
13574 negative economic conditions 0.230 0.378
14304 economic terminology 0.300 0.456
14747 supply‑related terms 0.234 0.382
Note: Generalized confidence intervals based on $k$-FWER control at $k=5$, $\alpha=5\%$. Descriptions generated using the method described in Appendix Section 6.2, using GPT-OSS 20B [OpenAI, 2025]. For bootstrapping, 10,00010,\!000 iterations are performed.
Once again, this procedure was cheap and fast to implement, and could be replicated and analyzed for sensitivity by any other researcher, quickly. Beyond filtering out greater than median activating dictionary features, no other choices were made to delimit the space of possible discoveries, or focus them in some way on the space of economic- or inflation-relevant concepts; the exact same models that yielded the results of Section 4.1 yielded the results in this section.
5 Conclusion
Existing literature in empirical economics and econometrics has long suggested the importance of open-ended discovery from high-dimensional or unstructured data. The framework proposed in this paper shows how new statistical procedures for high-dimensional multiple hypothesis testing, when combined with the latest innovations in interpretability methods for machine learning models, can facilitate open-ended, interpretable discovery at scale with higher practicality and higher fidelity than previously possible.
Importantly, the automaticity of the proposed framework makes it subject to very few researcher degrees of freedom, making it resilient to cherry-picking and motivated data-mining without compromising the purpose of discovery in the first place. This is especially true if one couples this framework with minimal preregistration efforts, e.g., simply publicly declaring what dictionary learning model, LLM, and automatic interpretation method are to be used for the analysis of a text, and if any of the features are to be filtered on principled grounds.
A reasonable possible criticism of the proposed framework is that the interpretations associated with dictionary learning methods are not credible. On this topic exactly, there has been an immense amount of attention from AI/ML research communities, which has, to date, primarily concluded that such dictionary features are useful and reliable for discovery, even if they do not uncover “atomic” or fundamental conceptual units in deep neural networks like LLMs [Leask et al., 2025, Peng et al., 2025]. The best practices and methods for moving from features to feature descriptions is an open and important question in the literature, and choices made on this front certainly affect interpretations of discoveries made by the proposed framework. However, as the empirical examples of the previous section make clear, straightforward automatic interpretation methods—which are completely data-driven, replicable, and not subject to motivated human reasoning—yield sensible outputs, scale well with the size of selected sets, can be further scrutinized by humans to assess their validity, and can be perturbed or modified cheaply to assess sensitivity to autointerpretation prompts. Future research related to this framework would seek to apply quantitative methods for validating feature descriptions at scale as well [Paulo et al., 2024, Movva et al., 2025].
The proposed framework is most naturally viewed as one tool of many in the empirical researcher’s toolkit for making discoveries from unstructured data. Using this framework alongside others that researchers may already be implementing is entirely complementary, and would only serve to deepen insights into possible interpretations of inference on unstructured data.

### v2

@@
 Abstract
-Social scientists are increasingly turning to unstructured datasets to unlock new empirical insights, e.g., estimating causal effects on text outcomes, measuring beliefs from open-ended survey responses.
-In such settings, unsupervised analysis is often of interest, in that the researcher does not want to pre-specify the objects of measurement or otherwise artificially delimit the space of measurable concepts; they are interested in “discovery.” This paper proposes a general and flexible framework for pursuing discovery from unstructured data in a statistically principled way.
-The framework leverages recent methods from the literature on machine learning interpretability to map unstructured data points to high-dimensional, sparse, and interpretable “dictionaries” of concepts; computes (test) statistics of these dictionary entries; and then performs selective inference on them using newly developed statistical procedures for high-dimensional exceedance control of the $k$-FWER under arbitrary dependence.
+Social scientists are increasingly turning to unstructured datasets to unlock new empirical insights, e.g., estimating descriptive statistics of or causal effects on quantitative measures derived from text, audio, or video data.
+In many such settings, unsupervised analysis is of primary interest, in that the researcher does not want to (or cannot) pre-specify all important aspects of the unstructured data to measure; they are interested in “discovery.” This paper proposes a general and flexible framework for pursuing discovery from unstructured data in a statistically principled way.
+The framework leverages recent methods from the literature on machine learning interpretability to map unstructured data points to high-dimensional, sparse, and interpretable “dictionaries” of concepts; computes statistics of dictionary entries for testing relevant concept-level hypotheses; performs selective inference on these hypotheses using algorithms validated by new results in high-dimensional central limit theory, producing a selected set (“discoveries”); and both generates and evaluates human-interpretable natural language descriptions of these discoveries.
 The proposed framework has few researcher degrees of freedom, is fully replicable, and is cheap to implement—both in terms of financial cost and researcher time.
@@
 1 Introduction
-Empowered by recent developments in machine learning and AI, researchers in the social sciences are increasingly leveraging sources of unstructured data—such as text, images, videos, and audio—in quantitative analyses.
-In economics, interest in unstructured data sources is especially widespread, ranging from using speech recordings from FOMC meetings to better understand monetary policy [Gorodnichenko et al., 2023]; to using videos of start-up pitches to study entrepreneurship and investment [Hu and Ma, 2025]; to using open-ended survey questions to probe the economic behavior and beliefs of individuals [Haaland et al., 2024]; to using mug shots to study judicial decisionmaking [Ludwig and Mullainathan, 2024]; to using visual art or written narratives to infer long-run living standards [Gorin et al., 2025, Lagakos et al., 2025]; to using interview transcripts to better understand the impacts of RCT treatments [Bergman et al., 2024, Krause et al., 2025].
-The promise of these unstructured data, fundamentally, is that they offer new opportunities to measure social phenomena that were previously unmeasurable.
-This includes both measures that a researcher is capable of pre-specifying as well as those that a researcher is not, and instead would like to “discover.” When modern AI/ML methods are brought to bear on these goals, the former falls under the heading of “supervised learning,” whereby the researcher typically makes predictions of their ex-ante known quantity of interest using the unstructured data as features, and the latter falls under the heading of “unsupervised learning,” which is less amenable to a prediction-based framework, and demands that AI/ML methods uncover latent structure in the unstructured data with minimal researcher input.
+Empowered by recent developments in machine learning and AI, researchers in the social sciences are increasingly leveraging sources of unstructured data—such as text, audio, images, and videos—in quantitative analyses.
+In economics, interest in unstructured data sources is especially widespread, ranging from using speech recordings from FOMC meetings to better understand monetary policy [Gorodnichenko et al., 2023]; to using videos of start-up pitches to study entrepreneurship and investment [Hu and Ma, 2025]; to using open-ended survey questions to probe the economic behavior and beliefs of individuals [Haaland et al., 2024]; to using mug shots to study judicial decisionmaking [Ludwig and Mullainathan, 2024]; to using visual art or written narratives to infer long-run living standards [Gorin et al., 2025, Lagakos et al., 2025]; to using qualitative interview transcripts to better understand the impacts of RCT treatments [Bergman et al., 2024, Krause et al., 2025].
+In the social sciences, the promise of these unstructured data, fundamentally, is that they offer new opportunities to measure social phenomena that were previously unmeasurable.
+This includes both measures that a researcher is capable of pre-specifying before seeing the data as well as those that a researcher is not, and instead would like to “discover” through the data itself.
+When modern AI/ML methods are brought to bear on these goals, the former typically falls under the heading of “supervised learning,” whereby the researcher makes predictions of their ex-ante known quantity of interest using the unstructured data as features, and the latter typically falls under the heading of “unsupervised learning,” which is less amenable to a prediction-based framework, and demands that AI/ML methods uncover latent structure in the unstructured data with minimal researcher input.
 Though new statistical and econometric frameworks have recently become available for interpretable and statistically principled analyses of unstructured data in the supervised learning setting (e.g., Angelopoulos et al.
 [2023], Ludwig et al.
-[2024], Carlson and Dell [2025]), it is an open question as to how best perform interpretable and statistically principled analyses of unstructured data in unsupervised learning settings that emphasize discovery.
+[2024], Carlson and Dell [2025]), it is an open question as to how one best performs interpretable and statistically principled analyses of unstructured data in unsupervised learning settings that emphasize discovery.
 The framework proposed in this paper provides one possible answer to this open question.
-Specifically, this paper proposes a general framework for conducting interpretable and rigorous statistical inference on unstructured data in the following way, starting from an unstructured dataset $\{Z_{i}\}_{i=1}^{n}$:
+Consider two motivating examples for the proposed framework, one focused on descriptive analysis, and another focused on causal inference.
+Both consider text as the unstructured data modality of interest, though the proposed framework applies more generally to other data modalities (see, e.g., the interpretability literature review in Section 2).
+Example 1 (Information treatment RCT (c.f., Bursztyn et al. [2023])).
+A researcher is interested in running a randomized controlled trial (RCT) with an information treatment.
+A key outcome of the experiment is the participant’s subjective reaction to the information treatment, and the researcher would like to discover any systematic differences in reactions across the treatment and control groups in an unsupervised way.
+As such, for all participants, the researcher collects text responses to open-ended questions eliciting reactions to the experimental intervention, and is interested in learning any relevant causal effects on the presence of concepts surfaced in these text responses.
+Example 2 (Open-ended survey analysis (c.f., Stantcheva [2024])).
+A researcher is interested in understanding the opinions and attitudes of some population of interest towards some particular (economic) policy.
+To learn about population attitudes without priming respondents to discuss particular topics, the researcher runs a survey on a representative sample from the population of interest that asks many open-ended questions about attitudes and beliefs, to which survey respondents reply in text.
+The researcher is interested in describing any prominent concepts that are surfaced in this open-ended survey.
+To achieve these motivating research goals—among many others—this paper proposes a general framework for conducting interpretable and rigorous statistical inference on unstructured data in the following way:
   1. 1.
-    Convert unstructured data to dictionary feature vectors: The researcher converts the unstructured data instances $Z_{i}\in\mathcal{Z}$ into high-dimensional, sparse, binary vectors $Y_{i}\in\{0,1\}^{p}$ using recently developed methods in dictionary learning for mechanistic interpretability of large language models (LLMs) and other deep neural networks (DNNs), i.e., $Y_{i}:=\mathtt{Dict}(Z_{i})$.
-Both theory and empirics corroborate that vector entries $Y_{ij}$ correspond to the presence of a monosemantic feature (i.e., a feature with a single, interpretable meaning) or “concept” $j$ in the given unstructured data instance $i$.
-There are $p\gg n$ dictionary features, e.g., we assume that $n$ may be on the order of $10^{2}$ or $10^{3}$, and $p$ is on the order of $10^{4}$ or $10^{5}$ (or larger).
-The breadth and depth of concepts spanned by these dictionary features is vast, as can be gleaned from the literature on dictionary learning for LLMs, and as will be observed concretely in the empirical applications of this paper.
-Intuitively, the dictionary should contain all possible (monosemantic) concepts that the LLM has learned for the purpose of next token prediction over a massive internet-scale corpus.
-However, due to sparsity imposed by design and suggested by machine learning theory, only a handful of these dictionary features will be present (“activate”) for any given unstructured data point.
+    The researcher begins with an unstructured dataset they are interested in analyzing.
+Each unstructured data point is mapped to a “feature dictionary,” which is a binary vector; each element of the vector indicates the presence (or lack thereof) of a single, human-interpretable (“monosemantic”) feature in that unstructured data point.
 [^1] For more discussion of what “monosemanticity” means and its importance for LLM interpretability, see, e.g., the discussion in Bricken et al.
 [2023].
+The creation of these interpretable feature dictionaries is accomplished using dictionary learning models newly popular in the machine learning interpretability literature for deep neural networks (DNNs) such as large language models (LLMs).
+These feature dictionary vectors are high-dimensional, and catalog a vast number of concepts.
+Intuitively, the feature dictionary should thought of as an inventory of all possible (monosemantic) concepts that, e.g., a LLM needed to learn for the purpose of next token prediction over a massive internet-scale corpus.
+By construction, these feature dictionary vectors are also sparse, i.e., only the few most “meaningful” concepts represented in the unstructured data point are cataloged as present by the dictionary.
+[^2] Meaningfulness, in a technical sense, relates to the reconstruction loss associated with learning the dictionary feature under a sparsity penalty; for more on this, see, e.g., Bricken et al.
+[2023].
+    Example 1 (Continued).
+    Each experimental participant’s post-treatment text response is passed through a LLM equipped with pretrained sparse autoencoder (SAE) [Bricken et al., 2023] (a popular implementation of a dictionary learning model); the SAE’s feature activations are pooled and thresholded to serve as a (binary) feature dictionary.
+There is one feature dictionary per text response (and one text response per experimental participant), each of which is of dimension on the order $p\approx 10^{5}$, though there are only on the order of $n\approx 10^{2}$ participants across both treatment and control groups.
+Feature number 1 in the dictionary may indicate the presence of sentences ending with exclamation points, feature number 2 may indicate the discussion of politics, and so on.
+[^3] For more examples of feature descriptions, see, e.g., https://www.neuronpedia.org/gemma-scope.
   2. 2.
-    Compute test statistics from the dictionary feature vectors: The researcher computes from this newly created dictionary feature indicator dataset $\{Y_{i}\}_{i=1}^{n}$ a vector of $p$ test statistics, $T_{n}\in\mathbb{R}^{p}$.
-These test statistics are intended to test $p$ hypotheses, one for each dictionary feature, and may be flexibly chosen by the researcher.
-Especially important and simple choices explored in this paper’s empirical examples include hypotheses testing means and differences-in-means, feature by feature.
+    The researcher, now with access to a new dataset composed of feature dictionaries, formulates $p$ feature-specific hypothesis tests and associated test statistics—one for each concept cataloged in the feature dictionary.
+The researcher has great flexibility in designing relevant, feature-level hypotheses and t-stats at this stage of the framework, which may employ data sources beyond the feature dictionaries themselves.
+    Example 1 (Continued).
+    The researcher computes a difference in means across treatment and control groups for each feature indicator (binary entry in the feature dictionary) to test a total of $p$ null hypotheses of zero average treatment effect (ATE) on the probability of a given feature appearing in a text response.
   3. 3.
-    Perform high-dimensional selective inference using the test statistics: Finally, the researcher invokes a high-dimensional selective inference procedure newly developed in this paper to control the $k$-FWER of the selected set of rejected hypotheses (“discoveries”).
-This procedure is asymptotically valid for small $k$, requires only very mild assumptions on the data generating process (which are met in the applications of interest), and is well-powered in the empirical examples considered.
-This procedure may also be of independent interest in other high-dimensional selective inference settings.
-The output of this framework is therefore a set of interpretable discoveries (rejections of dictionary feature derived tests) based on some set of unstructured data $\{Z_{i}\}_{i=1}^{n}$ with generalized familywise error rate guarantees, which may be relaxed or tightened as the researcher sees fit to permit more or fewer discoveries.
-Two notable use cases for this framework explored in the empirical applications of this paper include:
-  1. 1.
-    Estimating average treatment effects on dictionary features in RCTs, or testing, for treatment indicator $D_{i}$ and fixed probability of treatment $\pi$,
-    $H_{0,j}:E[Y_{ij}(1)-Y_{ij}(0)]=E[H_{i}Y_{ij}]=0,\quad H_{i}=\frac{D_{i}-\pi}{\pi(1-\pi)}.$
-  2. 2.
-    Estimating the probability of the presence of various concepts (measured by monosemantic dictionary features) in the population from which the unstructured dataset was sampled, or testing
-    $H_{0,j}:E[Y_{ij}]=0$
-    and then inverting these tests to form a “generalized” confidence set [Romano and Wolf, 2007].
+    All $p$ feature-specific hypotheses are tested using statistical procedures validated by novel results in high-dimensional multiple hypothesis testing and central limit theory.
+In particular, these procedures are shown to offer a novel formal guarantee over a generalized error rate for the selected set (the set of rejected null hypotheses) in high-dimensions, namely the $k$ familywise error rate ($k$-FWER) [Romano and Wolf, 2007]—the probability of making $k$ or more false rejections.
+This form of selective error control is less conservative that simple FWER control, [^4] Note that FWER is equivalent to $k$-FWER when $k=1$.
+leading to more powerful tests that permit the researcher to make more discoveries (rejections of true nulls).
+Moreover, the researcher may invert these tests to form generalized simultaneous confidence intervals, i.e., the probability that $k$ or more parameters are not contained by the random intervals is less than some desired level.
+These new formal results make only very mild assumptions on the data generating process, and may also be of independent interest in other high-dimensional selective inference settings.
+    Example 1 (Continued).
+    A fast, bootstrap-based algorithm is deployed to produce a critical value (or sequence of critical values) for rejecting each of the $p$ null hypotheses of no average treatment effect, guaranteeing $k$-FWER control in the selected set at level $\alpha$ (in large samples) for a researcher chosen value of $k$, even though $p\gg n$.
+These bootstrapped critical values may also be used to form generalized confidence intervals for the difference-in-means point estimates.
+  4. 4.
+    The output of the framework thus far is a selected set of rejected, feature-level null hypotheses (and corresponding generalized simultaneous confidence intervals and point estimates), a.k.a., “discoveries.” By virtue of the technology used to create them, however, the concepts encoded by the feature dictionary are only indexed numerically, and do not come with natural language descriptions—we know, e.g., that feature 1 likely encodes a human-interpretable concept, and that we rejected the null hypothesis associated with it, but we do not yet have an English description for what that concept is.
+As such, in the final stage of the framework, the researcher generates natural language descriptions of the concepts in the feature dictionary using a variant of a popular automatic interpretation (or “autointerp”) method that is localized to the unstructured dataset being explored.
+In order to rigorously evaluate the quality of these autointerp descriptions, we propose a new statistical formalization of a workhorse autointerp evaluation method, along with new estimators that permit statistically and conceptually principled autointerp quality evaluation.
+    Example 1 (Continued).
+    The researcher now has in hand some subset of the $p$ null hypotheses that have been rejected: a set of indices associated with particular elements of the feature dictionary.
+In order to provide a natural language description for each of the selected features for which there is evidence of a nonzero ATE (which are otherwise just identified by their numeric index), a dedicated “explainer LLM” reads examples of texts in which a given feature is present (or presents with high SAE activations) and offers an English interpretation for what concept is encoded by the feature; the credibility of this LLM-derived description of the feature is then evaluated with a held-out evaluation sample using newly proposed estimators.
+  5. 5.
+    The final output of the proposed framework is therefore: a set of feature-level discoveries (rejections of nulls) that come with generalized familywise error rate guarantees; natural (e.g., English) language descriptions of those discoveries; and numeric scores indicating the quality of those descriptions.
+    Example 1 (Continued).
+    The researcher discovers that the null of no average treatment effect is rejected for features indexed $\{4,101,5030,$...\}, meaning that with high probability the information treatment induces nonzero causal effects on the probability of the appearance of these features in texts collected across the treatment and control groups.
+Feature 4 is described by an explainer LLM as activating on “mentions of politics” and this description is evaluated to have a high “quality” score, i.e., the information treatment credibly causes increased discussion of politics; and so on.
 This particular framework for discovery from unstructured data conveys several scientific benefits.
 The first is that of having few researcher degrees of freedom.
-In many settings featuring exploratory analysis of unstructured data, there is concern that researchers may cherry pick aspects of their data to measure or cherry pick choices of measurements in order to reverse engineer a specific conclusion.
-By design, however, the framework proposed by this paper has very little room for motivated data mining, as it requires no “human-in-the-loop” to direct what quantities are ultimately measured.
-Off-the-shelf, open-source dictionary learning models generate $p$ features spanning a huge number of concepts that are well-cataloged and publicly documented, making it unlikely the researcher needs to define additional concepts to measure ad hoc, and making a failure to report hypotheses for certain features conspicuous.
-Moreover, prior to analysis, a researcher could simply preregister a choice of dictionary learning model, such as particular sparse autoencoder (SAE) for a particular LLM, limiting the possibility that a researcher applies this framework with many different dictionary learning methods to select on favorable results.
-The proposed framework is also inexpensive both in terms of financial cost and researcher burden; both empirical examples in this paper were computed on Google Colab notebooks using a single A100 GPU, access to which only costs tens of dollars per month.
-[^2] This pricing description is valid as of October 2025.
-Further, this framework is easy to implement, requiring no human-in-the-loop to label unstructured data instances, no additional AI/ML model training, and no additional data sources; the framework is nearly automatic in execution, and takes only minutes to run using the above-mentioned hardware.
-This automaticity also lends the framework to fast and easy replication, another important scientific benefit.
+In many settings featuring exploratory analysis of unstructured data, there is concern that researchers may cherry pick aspects of their data to measure in order to reverse engineer a specific conclusion.
+By design, however, the framework proposed by this paper has very little room for motivated data snooping, as it requires no “human-in-the-loop” to direct what quantities are ultimately measured or how those quantities are interpreted.
+Off-the-shelf, open-source dictionary learning models generate $p$ features spanning a huge number of concepts, making it implausible that the researcher needs to define additional concepts to measure ad hoc, and making a failure to report hypotheses for certain features conspicuous.
+Localized autointerp methods, especially when coupled with formal interpretation quality evaluation, further limit a researcher’s ability to guide analysis towards a preordained conclusion.
+Moreover, prior to analysis, a researcher can easily preregister a choice of dictionary learning model and explainer LLM, hampering the possibility that a researcher applies this framework with many different models to select on favorable results.
+[^5] Even such a method for pursuing motivated data snooping is unlikely to succeed, given that, e.g., SAE features tend to exhibit some degree of universality across architectures (c.f., Wang et al.
+[2025]).
+The proposed framework is also inexpensive both in terms of financial cost and researcher burden; both empirical examples in this paper were computed on Google Colab notebooks using a single A100 GPU, access to which only costs tens of dollars per month (and is free for students).
+[^6] This pricing description is valid as of December 2025.
+This framework is easy to implement, requiring no human-in-the-loop to label unstructured data points, no additional AI/ML model training, and no additional data sources not already possessed by the researcher; the framework is nearly automatic in execution, and takes on the order of a single hour to run using the above-mentioned hardware (and with access to LLMs via API, it may only take minutes).
+This automaticity also lends the framework to fast and easy replication and sensitivity analysis, other important scientific benefits.
 Thus, the primary contributions of this paper are:
@@
   2. 2.
-    The development of novel high-dimensional multiple hypothesis testing procedures for controlling $k$-FWER exceedance, facilitated by extensions of high-dimensional central limit theory.
-The remainder of this paper is structured as follows: Section 2 discusses related literature; Section 3 describes the framework in detail; Section 4 applies the framework to two recent papers making use of unstructured data for discovery [Bursztyn et al., 2023, Stantcheva, 2024]; and Section 5 concludes.
+    Novel, general-purpose results in high-dimensional central limit theory and high-dimensional multiple hypothesis testing that enable statistical procedures for controlling $k$-FWER in high-dimensional settings.
+  3. 3.
+    A new conceptual and statistical formalization of a popular framework for autointerp scoring, including novel estimators for implementing it.
+The remainder of this paper is structured as follows: Section 2 discusses related literature; Section 3 describes the framework in detail (applied researchers are primarily encouraged to read Sections 3.1, 3.4, and 3.5); Section 4 applies the framework to two recent papers making use of unstructured data for discovery [Bursztyn et al., 2023, Stantcheva, 2024]; and Section 5 concludes.
 2 Related Literature
@@
 Motivated by the capabilities of modern AI/ML methods for learning from unstructured economic data (see, e.g., Mullainathan and Spiess [2017], Gentzkow et al.
-[2019a], Dell [2025]), new econometric and statistical frameworks have been developed to facilitate principled statistical inference on low-dimensional features (predictions) learned from unstructured datasets under supervision, e.g., Angelopoulos et al.
+[2019a], Ash and Hansen [2023], Dell [2025]), new econometric and statistical frameworks have been developed to facilitate principled statistical inference on low-dimensional features (predictions) learned from unstructured datasets under supervision, e.g., Angelopoulos et al.
 [2023], Ludwig et al.
@@
 Modarressi et al.
-[2025] uses LLM reasoning as a tool for dimensionality reduction on the space of concepts present in text (i.e., as a low-dimensional hypothesis generation methodology) and incorporates human validation of LLM-selected hypotheses, with estimation implemented via sample splitting; in contrast, the framework proposed in this paper leverages recent advances in interpretability methods for LLMs in order to directly analyze the high-dimensional space of concepts present in text, and uses state-of-the-art autointerpretation methods to describe discovered concepts without further researcher intervention, and without requiring data splitting.
+[2025] uses LLM reasoning as a tool for dimensionality reduction on the space of concepts present in text (i.e., as a low-dimensional hypothesis generation methodology) and incorporates human validation of LLM-selected hypotheses.
+In contrast, the framework proposed in this paper leverages recent advances in interpretability methods for LLMs in order to directly analyze the high-dimensional space of concepts present in text, and uses state-of-the-art autointerpretation methods to describe discovered concepts without further researcher intervention, similarly evaluating these descriptions using quantitative methods that require no researcher discretion.
 As such, these frameworks offer different profiles of scientific benefits and researcher costs for the problem of causal inference on text.
@@
 However, the proposed framework tackles the high-dimensional discovery problem using a novel high-dimensional selective inference procedure that maintains a high degree of interpretability, whereas Ludwig et al.
-[2017] modifies the target of discovery to improve power, trading off interpretability to do so.
+[2017] modify the target of discovery to improve power, trading off interpretability to do so.
 High-dimensional selective inference.
@@
 The methods in this literature span asymptotic and finite sample valid frameworks, and low- and high-dimensional settings.
-In particular, for the problem considered in this paper, the literature on asymptotically valid, high-dimensional selective inference is most relevant, as such methods permit making very few assumptions about the data (c.f., assuming independence or PRDS of the p-values in the case of Benjamini and Hochberg [1995]) and readily incorporate resampling methods that improve power (relative to guarding against worst-case dependence, c.f., Bonferroni corrections or the method of Benjamini and Yekutieli [2001]).
-These new procedures are built by extending recent work on high-dimensional central limit theory for order statistics [Chernozhuokov et al., 2022, Ding et al., 2025].
-The proposed statistical procedures are the first stated of their kind that enable asymptotic high-dimensional exceedance control of the $k$-FWER for test statistics that may be arbitrarily correlated with one another, complementing existing methods for asymptotic high-dimensional control of the FWER [Belloni et al., 2018] and the FDR under weak dependence [Liu and Shao, 2014, Belloni et al., 2018].
+In particular, for the problem considered in this paper, the literature on asymptotically valid, high-dimensional selective inference is most relevant, as such methods permit making very few assumptions about the data (c.f., assuming independence or PRDS of p-values in the case of Benjamini and Hochberg [1995]) and readily incorporate resampling methods that improve power (relative to guarding against worst-case dependence, c.f., Bonferroni corrections or the method of Benjamini and Yekutieli [2001]).
+The new results for high-dimensional multiple hypothesis testing appearing in this paper are derived by extending recent work on high-dimensional central limit theory for order statistics [Chernozhuokov et al., 2022, Ding et al., 2025].
+The formal results presented in this paper are the first stated of their kind that enable asymptotic high-dimensional control of the $k$-FWER for test statistics that may be arbitrarily correlated with one another, complementing existing methods for asymptotic high-dimensional control of the FWER [Belloni et al., 2018] and the FDR under weak dependence [Liu and Shao, 2014, Belloni et al., 2018].
 Hypothesis generation.
@@
 From the perspective of this literature, the proposed framework may be viewed as leveraging dictionary learning as a high-dimensional hypothesis generator, on top of which discovery is performed at the “scientific” stage of inquiry using high-dimensional multiple hypothesis testing procedures.
-As such, the framework in this paper contributes to the literature on hypothesis generation by providing a statistically principled pathway from generation to evaluation—one which involves little to no researcher discretion about what generated hypotheses should be tested.
+As such, the framework in this paper contributes to the literature on hypothesis generation by providing a statistically principled pathway from hypothesis generation to hypothesis evaluation given access to only one (potentially small) unstructured dataset—a pathway that involves little to no researcher discretion about what generated hypotheses should be tested.
 The recent framework of Movva et al.
@@
 Interpretability for LLMs and other DNNs.
-There is a large literature in computer science on machine learning interpretability (see, e.g., Doshi-Velez and Kim [2017] for a foundational agenda).
+There is a large literature in computer science on machine learning interpretability (see, e.g., Doshi-Velez and Kim [2017]).
 The machine learning interpretability methods leveraged by this framework originate from a nascent though highly active literature known as “mechanistic interpretability,” which seeks to develop methods for interpreting LLM behavior via quantitative analyses of model internals (e.g., activations, weights).
@@
 [2024], i.e., sparse autoencoders (SAEs).
-SAEs are autoencoders attached to the residual streams (or other model internals) of a LLM, with hidden layers being many orders larger in dimension than the residual stream, and which are trained with sparsity-inducing penalization on reconstruction loss.
+SAEs are very wide autoencoders inserted at various layers of an LLM that are trained to intercept and reconstruct activations from model internals (such as the residual stream, MLPs or the like) under a sparsity-inducing penalization on a reconstruction loss.
 Under the “linear representation hypothesis” and “superposition hypothesis” (see, e.g., Bricken et al.
-[2023] for more discussion), these sparse autoencoders are thought to act as an overcomplete basis of the space of concepts in text leveraged by a LLM to make next token predictions, encouraging learned features to be monosemantic.
+[2023] for more discussion), these sparse autoencoders are thought to act as an overcomplete basis of the space of concepts in text leveraged by a LLM to make next token predictions, encouraging learned autoencoder latents to have monosemantic interpretations.
 There has been much debate about the success of SAEs (e.g., Leask et al.
@@
 SAEs and other dictionary learning methods generate features, but not natural language feature descriptions.
-Based on the popularity of these methods, a related literature on “autointerpretability” methods has become active, which seeks to coherently use LLMs to describe the features discovered by dictionary learning methods at scale.
+Based on the popularity of these methods, a related literature on “automatic interpretability” (or “autointerp”) methods has become active, which seeks to coherently use LLMs to describe the features discovered by dictionary learning methods at scale.
 Recent important papers in this literature include: Bills et al.
@@
 [2024].
-The framework proposed in this paper primarily uses insights from Paulo et al.
-[2024] to create “local interpretations” that are most relevant to the data generating process being considered, and otherwise leverages autointerpetations from Lieberum et al.
-[2024].
-Though the empirical examples considered in this paper handle one of the most popular unstructured date types—text—for which dictionary learning-based interpretability methods are the most well-developed, dictionary learning techniques have been successfully applied to many other modalities, including audio and images [Abdulaal et al., 2024, Bhalla et al., 2024, Fry, 2024, Daujotas, 2024, Pluth et al., 2025].
+Specifically, the autointerp framework proposed in this paper adapts and formalizes insights from the “detection scoring” method of Paulo et al.
+[2024] in a way that emphasizes “locality”: maintaining a high degree of relevance to the unstructured data generating process being considered.
+Though the empirical examples considered in this paper handle one of the most popular unstructured data types—text—for which dictionary learning-based interpretability methods are the most mature, dictionary learning techniques have been successfully applied to many other modalities, including audio and images [Abdulaal et al., 2024, Bhalla et al., 2024, Fry, 2024, Daujotas, 2024, Pluth et al., 2025].
+Future work would seek to apply the proposed framework to unstructured datasets in these other data modalities.
 3 Framework
 3.1 Setup
-Let $n]:=\{1,\dots,n\}$, and let $x\mapsto\log x$ be the natural logarithm.
-The researcher has access to a dataset of size $n$, $\{(W_{i},Z_{i})\}_{i=1}^{n}$, which is sampled i.i.d.
-from some (super-)population of interest $P$.
-The $Z_{i}\in\mathcal{Z}$ are unstructured data instances and the $W_{i}\in\mathcal{W}$ are any other observed covariates of interest (e.g., for one empirical example considered in the following section, $W_{i}=D_{i}\in\mathcal{W}=\{0,1\}$, a binary treatment status indicator).
+By way of notation, define $n]:=\{1,\dots,n\}$, and let $x\mapsto\log x$ be the natural logarithm.
+Consider that the researcher has access to a dataset of size $n$, $\{(W_{i},Z_{i})\}_{i=1}^{n}$, which is sampled i.i.d.
+from some (super-)population of interest represented by distribution $P$.
+The $Z_{i}\in\mathcal{Z}$ are unstructured data instances and the $W_{i}\in\mathcal{W}$ are any other observed covariates of interest.
 The space $\mathcal{Z}$ is typically high-dimensional and semantically poor, e.g., if each $Z_{i}$ was a $244\times 244$ pixel image, $\mathcal{Z}$ might then be the space of all $3\times 244\times 244$ arrays of RGB values, or if $Z_{i}$ was text that was truncated up to some maximum length, $\mathcal{Z}$ might be the space of all binary matrices of a certain dimensionality, which are concatenated one-hot encodings for each word, term, or token in the text with respect to a specific pre-defined vocabulary.
-The researcher has access to the function $\texttt{Dict}:\mathcal{Z}\to\{0,1\}^{p}$, which maps an unstructured data observation to a a sparse binary vector.
-This function is computed by passing an unstructured data point to a DNN equipped with a dictionary learning model (e.g., a LLM equipped with pretrained SAE when $Z_{i}$ are texts), recording the feature activations from the dictionary model, and then aggregating and binarizing these dictionary feature activations to form a single indicator for each feature in the dictionary $j\in[p]$.
+Example 1 (Continued).
+The $Z_{i}$ is a text response elicited for experimental participant $i$; the $W_{i}\in\mathcal{W}=\{0,1\}$ is a treatment status indicator for experimental participant $i$.
+The researcher has access to a function which creates feature dictionaries from unstructured data points, $\texttt{Dict}:\mathcal{Z}\to\{0,1\}^{p}$.
+This function is computed by passing an unstructured data point to a DNN equipped with a dictionary learning model (e.g., a LLM equipped with pretrained SAE when $Z_{i}$ are texts); recording the feature activations from the dictionary model; and then pooling and binarizing these dictionary feature activations to form a single indicator for each feature in the dictionary, indexed $j\in[p]$.
 There is a large possible space of Dict functions the researcher could implement, though an especially straightforward choice for the use case of text data is implementing a Dict function based on the SAE activations at a single layer of the DNN, for which the $j$-th entry of the output is equal to 1 if the $j$-th dictionary feature activated on any token of the input text, and 0 otherwise:
 […]
-We denote $Y_{i}:=\mathtt{Dict}(Z_{i})\in\{0,1\}^{p}$ as the output of the dictionary transformation of the unstructured data.
-As such, each $Y_{ij}\in\{0,1\}$ has the natural interpretation that a particular dictionary feature $j$ activated for the unstructured data instance $i$, and, e.g., $\bar{Y}_{\bullet,j}:=n^{-1}\sum_{i=1}^{n}Y_{ij}$ corresponds to an estimate of the probability of the presence of a feature $j$ in the population $P$ of interest.
+We denote $Y_{i}:=\mathtt{Dict}(Z_{i})\in\{0,1\}^{p}$ as the output of the dictionary transformation of the unstructured data: the “feature dictionary.” As such, each “feature indicator” or “dictionary entry” $Y_{ij}\in\{0,1\}$ has the natural interpretation that a particular dictionary feature $j$ activated for the unstructured data instance ii.
 [^7] Dictionary feature activations are typically positively valued scalars, where magnitude is thought to correspond to some notion of intensity of activation; other works in mechanistic interpretability consider using these raw activation values directly, perhaps with max or average pooling across tokens.
 However, presently, magnitude of activation is not a well-understood or highly interpretable quantitative property of dictionary learning methods, and as such the default Dict function advocated for in the proposed framework does not incorporate this information.
-That said, researchers interested in extending this framework may find value in alternative Dict functions that leverage activation intensity.
-The crux of the proposed framework is that inference on functionals of $Y_{ij}$ is desirable because:
+That said, leveraging feature activation intensity is empirically shown to be useful for deriving good autointerpretation generation strategies; see Section 3.4 for more on this.
+Example 1 (Continued).
+Each text response $Z_{i}$ is passed through an open-source LLM equipped with pretrained SAE.
+The SAE activations for each token of $Z_{i}$ at the $l$-th layer of the LLM are max-pooled and then thresholded at zero, yielding a $p\approx 10^{5}$-dimensional binary vector $Y_{i}$ for each experimental participant $i$.
+Each $Y_{ij}\in\{0,1\}$ is a binary indicator of a particular concept being expressed in experimental participant $i$’s text response, e.g., $Y_{i1}$ could indicate whether or not experimental participant $i$’s text response contains phrases indicating uncertainty.
+The crux of the proposed framework is that inference on functionals of (the distribution of) each $Y_{ij}$ is desirable because:
   1. 1.
-    Each $Y_{ij}$ is monosemantic, and therefore functionals of $Y_{ij}$ for a given $j$ are interpretable, e.g., $E_{P}[Y_{ij}]$ is the probability of monosemantic feature $j$ appearing in an unstructured data observation in population $P$.
+    Each $Y_{ij}$ is monosemantic (indicates the presence of a single, intelligible concept) and therefore functionals for $Y_{ij}$ for any given $j$ (and associated hypotheses) are interpretable, e.g., $E_{P}[Y_{ij}]$ is the probability of human-interpretable feature $j$ appearing in an unstructured data observation in population $P$.
   2. 2.
-    Collectively, the $p$ features in $Y_{ij}$ for any given $i$ are exhaustive of some vast space of concepts of interest to the researcher; intuitively, they are the set of all monosemantic features that a LLM needed to learn to perform well on next token prediction for a massive, internet-scale corpus of text (or other similarly massive unstructured data source).
-This means the researcher need not pre-specify any particular concepts of interest, reducing researcher degrees of freedom without limiting discovery.
+    Collectively, the $p$ features in $Y_{i}$ for any given $i$ are exhaustive of some vast space of concepts of interest to the researcher.
+Heuristically, they encode the set of all concepts that a LLM needed to learn to perform well on next token prediction for a massive, internet-scale corpus of text (or other similarly massive unstructured data source).
+This means the researcher need not specify ex-ante or ex-post any particular concepts of interest, reducing the conceptual and statistical complications associated with data snooping without limiting discovery.
 However, inference on functionals of $Y_{ij}$ for each $j$ is also challenging, because:
@@
     Each $Y_{i}$ is sparse, and the intent of the analysis is discovery, so the desired form of selective (familywise) error control cannot be too conservative.
-What is needed, then, for principled inference on functionals of $Y_{ij}$ is a high-dimensional selective inference procedure with control over a generalized error rate.
+What is needed, then, for principled inference on functionals of $Y_{ij}$ is a high-dimensional selective inference procedure with control over a generalized selective error rate.
 To be as well-powered as possible, we want to focus on testing procedures that employ resampling methods for estimating the true covariance across test statistics, such that conservative protection against worst-case dependence is not required.
-To keep things as general as possible, we also want to allow for only asymptotically valid test statistics and p-values, as is common to much of econometric analysis.
+To keep things as general as possible, we also want to allow for only asymptotically valid test statistics and p-values, as is common in much of econometric analysis.
 In the following sections, theory and corresponding statistical procedures are developed to achieve exactly these aims, and are stated as generally as possible to accommodate settings of interest even beyond analysis of unstructured data.
-To pursue inference, consider defining $p\gg n$ “one-sided” hypotheses $\{H_{0,j}\}_{j\in[p]}$, where
-$H_{0,j}:\theta_{j}(P):=E_{P}[X_{ij}]\leq 0,\quad X_{ij}=h(W_{i},Y_{ij}),$
-for some measurable function $h$.
-Two choices of $h$ considered in the empirical examples in this paper include $h(W_{i},Y_{ij})=Y_{ij}$ (testing probabilities of dictionary feature activation) and $h(W_{i},Y_{ij})=\frac{D_{i}-\pi}{\pi(1-\pi)}Y_{ij}$ (testing differences in probabilities of activation across groups under known, fixed group assignment), though many others are possible, lending to the flexibility of the framework.
-As shorthand, we will denote $X:=\{X_{i}\}_{i=1}^{n}$.
-Consider also defining a set of “two-sided” hypotheses of interest $\{\tilde{H}_{0,j}\}_{j\in[p]}$ as
-$\tilde{H}_{0,j}:\theta_{j}(P):=E_{P}[X_{ij}]=0.$
+To set up multiple hypothesis testing, consider defining $p\gg n$ “one-sided” feature/concept-level hypotheses $\{\tilde{H}_{0,j}\}_{j\in[p]}$, where
+$\tilde{H}_{0,j}:\theta_{j}(P):=E_{P}[X_{ij}]\leq 0,\quad X_{ij}=h(W_{i},Y_{ij}),$
+for some measurable function $h$. Consider also defining a set of “two-sided” hypotheses of interest $\{H_{0,j}\}_{j\in[p]}$ as
+$H_{0,j}:\theta_{j}(P):=E_{P}[X_{ij}]=0.$
+As shorthand, we will denote $X:=\{X_{i}\}_{i=1}^{n}$. Many useful choices of $h$ are possible under this framework, lending to its flexibility.
+Example 1 (Continued).
+To infer average treatment effects on dictionary features, for treatment indicator $W_{i}$, fixed probability of treatment $\pi$ (bounded away from zero and one), and feature indicator potential outcomes $Y_{ij}(w)$ under SUTVA,
+$H_{0,j}:E[Y_{ij}(1)-Y_{ij}(0)]=E[X_{ij}]=E\left[\frac{W_{i}-\pi}{\pi(1-\pi)}Y_{ij}\right]=0.$
+In other words, we set $X_{ij}=h(W_{i},Y_{ij}):=\frac{W_{i}-\pi}{\pi(1-\pi)}Y_{ij}$, a Horvitz–Thompson transformation for computing a difference in means.
 We may form test statistics for such hypotheses as
-$T_{n,j}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{ij},\quad T_{n}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{i}.$
-For the purposes of developing high-dimensional central limit theory later on, further define
+$T_{n,j}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{ij},\quad T_{n}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{i},$
+(and taking absolute values, as appropriate). For the purposes of developing high-dimensional central limit theory later on, further define
 $S_{n,j}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{ij}-E[X_{ij}]),\quad S_{n}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-E[X_{i}]).$
@@
 data.
-Further, for any subset of indices $K\subseteq[p]$, we will denote $S_{n,K}:=(S_{n,j}:j\in K)^{\mathtt{T}}$.
-3.2 High-Dimensional $k$-FWER Exceedance Control
-We construct methods that provide high-dimensional $k$-FWER exceedance control asymptotically, or control of the probability of making $k$ or more rejections of true nulls in the large sample limit when $p$ is growing (much) faster than $n$ in the relevant asymptotic thought experiment.
-Naturally, $k$-FWER control for $k=1$ is control of the FWER, and larger choices of $k$ permit more discoveries.
+Further, for any subset of indices $K\subseteq[p]$, we will denote the subvector $S_{n,K}:=(S_{n,j}:j\in K)^{\mathtt{T}}$.
+3.2 High-Dimensional $k$-FWER Control
+We construct methods that provide high-dimensional $k$-FWER control asymptotically, i.e., control of the probability of making $k$ or more rejections of true nulls in the large sample limit when $p$ is growing (much) faster than $n$ in the relevant asymptotic thought experiment.
+Naturally, $k$-FWER control for $k=1$ is control of the FWER, and larger choices of $k$ make the permit more discoveries.
 To proceed, we will adapt strategies from Romano and Wolf [2007] for achieving large sample $k$-FWER control to the high-dimensional setting.
 The testing procedures of Romano and Wolf [2007] rely on the asymptotic validity of bootstrap approximations of the distribution of the $k$-th largest coordinate (or “$k$-max”) of $S_{n}$.
-Thus, a bootstrap procedure for the $k$-max statistic that is asymptotically valid in high-dimensions would allow one to run the algorithms of Romano and Wolf [2007] in the high-dimensional setting with their stated control under minimal modifications.
-High-dimensional bootstraps must be supported by appropriate high-dimensional central limit theory.
+Thus, the core intuition for results stated in this section is that, by building on a bootstrap procedure for the $k$-max statistic that is asymptotically valid in high-dimensions, we may salvage the general construction of the algorithms of Romano and Wolf [2007] to achieve analogous generalized selective error guarantees in high-dimensions to the low-dimensional setting originally considered by them.
+High-dimensional bootstraps must be validated by appropriate high-dimensional central limit theory.
 Fortunately, the recent work of Ding et al.
@@
 As such, for i.i.d.
-data, the case considered in this paper, we may simply require for some fixed $B<\infty$ and fixed $b_{1},b_{2}$ with $b_{1}\leq b_{2}$: (i) $E\left[\exp\left(\left|X_{ij}\right|/B\right)\right]\leq 2$; (ii) $b_{1}^{2}\leq E[X_{ij}^{2}]$; and (iii) $E[X_{ij}^{4}]\leq B^{2}b_{2}^{2}$.
-Note that, in many use cases of interest for dictionary learned features, the $X_{ij}$ are both bounded and studentized, aiding the plausibility of these assumptions.
-In practice, some fraction of dictionary features may be “dead” for any given dataset, meaning they are degenerate at zero [Bricken et al., 2023].
-Of course, these features are not of interest for the purposes of discovery.
-This phenomenon is tolerated by the framework so long as the count of degenerate features is smaller than any relevant pkp-k, in which case, in any finite sample, the $k$-th largest coordinate of all relevant statistics is invariant to whether or not the degenerate coordinates are dropped.
-As such, conceptually, one should think of $j$ as ranging over the $p$ non-degenerate features.
-In practice, to perform studentization, it is recommended that these degenerate coordinates are explicitly filtered out; studentization is discussed further in Section 3.3.
+data, the case considered in the applications of this framework, we may simply require for some fixed $B<\infty$ and fixed $b_{1},b_{2}$ with $b_{1}\leq b_{2}$: (i) $E\left[\exp\left(\left|X_{ij}\right|/B\right)\right]\leq 2$; (ii) $b_{1}^{2}\leq E[X_{ij}^{2}]$; and (iii) $E[X_{ij}^{4}]\leq B^{2}b_{2}^{2}$.
+Note that, in many use cases of interest for dictionary learned features, the $X_{ij}$ will be both bounded (almost surely) and studentized, making these assumptions highly plausible.
+[…]
 Assumption R.
 Assume that $B_{n}^{2}\log^{5}(pn)=o(n)$.
-This is the key rate condition needed for (sup-norm) Gaussian or bootstrap approximation error to go to zero asymptotically under the high-dimensional CLTs discussed in both Ding et al.
+This is the key rate condition needed for (sup-norm) Gaussian or bootstrap approximation error to go to zero asymptotically under the high-dimensional CLTs discussed in both the works of Ding et al.
 [2025] and Chernozhuokov et al.
@@
 That this rate condition permits $p\gg n$ is an important positive result in high-dimensional central limit theory [Chernozhukov et al., 2017, 2023].
+Both Assumptions M and R are the state of the art in the high-dimensional central theory from which this framework draws [Chernozhuokov et al., 2022, Ding et al., 2025].
+Indeed, these conditions are the exact same as those used in both Chernozhuokov et al.
+[2022] and Ding et al.
+[2025].
+Notably, these assumptions do not require non-degeneracy of $\Sigma$, a well-established and beneficial feature of these results [Chernozhukov et al., 2023].
+Under assumptions of non-degenerate $\Sigma$, improved rate conditions may be used in place of Assumption R.
 We will focus on the Gaussian multiplier bootstrap as our high-dimensional bootstrap method. We define the Gaussian multiplier bootstrap quantity
@@
 where $\xi_{i}\overset{iid}{\sim}N(0,1)$ and $\bar{X}_{n}:=n^{-1}\sum_{i=1}^{n}X_{i}$.
-We also introduce $x\mapsto x_{[k]}$ as notation for the function that selects the $k$-th largest coordinate of a vector $x$, i.e., $x_{[k]}=k\text{-}\!\max(x)$ in the notation of Romano and Wolf [2007].
+We also introduce $x\mapsto x_{[k]}$ as notation for the function that selects the $k$-th largest coordinate of a vector $x$, and use notation $x\mapsto k\text{-}\!\max(x)$ to denote the $k$-the largest element of a set, per the notation of Romano and Wolf [2007].
 For the purposes of bootstrapping, we will denote the data-conditional probability measure $P^{B}(\cdot):=P(\cdot\mid X)$.
-With these assumptions and notations in place, we now introduce statistical procedures for controlling the $k$-FWER for small $k$, based on the step-wise algorithms of Romano and Wolf [2007].
+With these assumptions and notations in place, we now introduce statistical procedures for controlling the $k$-FWER for small $k$ in large samples, based on the step-wise algorithms of Romano and Wolf [2007].
 {restatable}
-[High-dimensional $k$-FWER exceedance control for small $k$, one-sided]theoremSmallk Consider the method of Algorithm 2.1 or 2.2 in Romano and Wolf [2007] with test statistics $T_{n}$ of hypotheses $\{H_{0,j}\}_{j\in[p]}$ and critical values $\hat{c}_{n,K}(1-\alpha,k)$ given by the $1-\alpha$ quantile of $S_{n,K,[k]}^{B}$ under $P^{B}$.
+[High-dimensional $k$-FWER control for small $k$, one-sided]theoremSmallk Consider the method of Algorithm 2.1 or 2.2 in Romano and Wolf [2007] with test statistics $T_{n}$ of hypotheses $\{\tilde{H}_{0,j}\}_{j\in[p]}$ and critical values $\hat{c}_{n,K}(1-\alpha,k)$ given by the $1-\alpha$ quantile of $S_{n,K,[k]}^{B}$ under $P^{B}$.
 Assume that $k$ is fixed (i.e., not growing with $n,p$).
@@
   2. (ii)
-    If $H_{0,j}$ is false and $\theta_{j}(P)\gg B_{n}\sqrt{\log p/n}$, then the probability that the step-down method rejects $H_{0,j}$ tends to 1.
+    If $\tilde{H}_{0,j}$ is false and $\theta_{j}(P)\gg B_{n}\sqrt{\log p/n}$, then the probability that the step-down method rejects $\tilde{H}_{0,j}$ tends to 1.
+This result is the natural high-dimensional analog of Theorem 3.1 (as well as Theorem 3.3) of Romano and Wolf [2007].
+It says that, so long as $k$ is small (in the relevant asymptotic sense), Algorithm 2.1 or Algorithm 2.2 of Romano and Wolf [2007] may be used to asymptotically achieve control of the $k$-FWER at any desired level $\alpha$ in high-dimensional settings, and moreover that these algorithms are consistent against local alternatives that are not shrinking towards zero too quickly [^10] Under Assumption R, note that $B_{n}\sqrt{\log p/n}=o(1)$.
+(i.e., intuitively, in sufficiently large samples, we are able to appropriately reject a false null so long as $\theta_{j}(P)=E_{P}[X_{ij}]$ is not very close to zero).
+The proof of this result is available in Appendix Section 6.4, and, as previously alluded to, relies on the insight that the low-dimensional $k$-FWER controlling algorithms of Romano and Wolf [2007] can be adapted to and analyzed in the high-dimensional setting using new results in high-dimensional central limit theory for order statistics (e.g., Ding et al.
+[2025]).
+In the following theorem, we state the sibling procedure for high-dimensional control of the $k$-FWER in two-sided testing settings; it is the high-dimensional analog of Theorem 3.2 of Romano and Wolf [2007].
 {restatable}
-[High-dimensional $k$-FWER exceedance control for small $k$, two-sided]theoremSmallktwo Consider the method of Algorithm 2.1 or 2.2 in Romano and Wolf [2007] with test statistics $|T_{n}|$ of hypotheses $\{\tilde{H}_{0,j}\}_{j\in[p]}$ and critical values $\hat{c}_{n,K}(1-\alpha,k)$ given by the $1-\alpha$ quantile of $|S_{n,K}^{B}|_{[k]}$ under $P^{B}$.
+[High-dimensional $k$-FWER control for small $k$, two-sided]theoremSmallktwo Consider the method of Algorithm 2.1 or 2.2 in Romano and Wolf [2007] with test statistics $|T_{n}|$ of hypotheses $\{H_{0,j}\}_{j\in[p]}$ and critical values $\hat{c}_{n,K}(1-\alpha,k)$ given by the $1-\alpha$ quantile of $|S_{n,K}^{B}|_{[k]}$ under $P^{B}$.
 Assume that $k$ is fixed (i.e., not growing with $n,p$).
@@
   2. (ii)
-    If $\tilde{H}_{0,j}$ is false and $|\theta_{j}(P)|\gg B_{n}\sqrt{\log p/n}$, then the probability that the step-down method rejects $\tilde{H}_{0,j}$ tends to 1.
-Importantly, the procedures of Theorems R and R are only valid for small $k$.
-To see this explicitly, note that the sup-norm bootstrap or Gaussian approximation error in the high-dimensional CLTs for the $k$-th largest coordinate introduced in Ding et al.
+    If $H_{0,j}$ is false and $|\theta_{j}(P)|\gg B_{n}\sqrt{\log p/n}$, then the probability that the step-down method rejects $H_{0,j}$ tends to 1.
+To comment explicitly on why the procedures of Theorems R and R are only valid for small $k$, note that the sup-norm bootstrap or Gaussian approximation error in the high-dimensional CLTs for the $k$-th largest coordinate introduced in Ding et al.
 [2025] only goes to zero if
@@
 As such, the theory of Ding et al.
-[2025] is unsuitable for making progress on FDP exceedance control, which requires a $k$ that may grow linearly with $p$.
-An important goal of future research would be to establish valid high-dimensional FDP exceedance control using appropriate high-dimensional central limit theory.
+[2025] is unsuitable for making progress on FDP exceedance control, which requires a $k$ that may grow linearly (“quickly”) with $p$.
+A useful goal of future research would be to establish valid high-dimensional FDP exceedance control using further extensions of high-dimensional central limit theory.
+However, as will be observed in the empirical applications of this paper, small $k>1$ are often sufficient to power meaningful discoveries in relevant social science settings at conventional choices of $\alpha$.
+[^11] For applied researchers looking for guidance on choice of $k$, note that the given approximation error rate suggests a choice of $k$ on the order of $\log(pn)$.
 3.3 High-Dimensional CLT for $k$-max of Approximate Means
-In practice, hypothesis testing with studentized statistics will be important for powering discoveries beyond just frequently occurring dictionary features (which are, e.g., high variance features for Bernoulli random variables when $X_{ij}=Y_{ij}\in\{0,1\}$).
-To facilitate studentization, we will need a high-dimensional CLT that permits statistics with small estimation errors.
-We therefore generalize the high-dimensional CLTs of Ding et al.
+Multiple hypothesis testing with studentized statistics is often important for the purposes of “balance”: that all hypothesis tests considered are similarly powered, contributing to selective error control similarly.
+For example, when $X_{ij}=Y_{ij}\in\{0,1\}$, more frequently occurring dictionary features are higher variance features, and we are typically interested in powering discoveries beyond just these frequently occurring concepts.
+To facilitate studentization, we require a high-dimensional CLT that considers statistics with small estimation errors, as studentization relies on estimated variances.
+We therefore generalize the high-dimensional central limit theorems of Ding et al.
 [2025] to handle “approximate means” in the parlance of Belloni et al.
-[2018], stated in the following lemmas.
+[2018], stated in the following lemma.
 {restatable}
@@
 $\sup_{t\in\mathbb{R}}\left|P\left(\hat{S}_{n,[k]}\leq t\right)-P\left(N(0,\Sigma)_{[k]}\leq t\right)\right|\to 0.$
+This lemma is a natural generalization of Lemma A.6 of Ding et al.
+[2025], which proves the special case for $R_{n}=0$.
+This lemma then indicates that, for appropriately small $k$, in the large sample limit, the distribution of the $k$-th largest coordinate of a scaled sum of independent random vectors is (uniformly) well-approximated by the distribution of the $k$-th largest coordinate of a Gaussian random vector with variance matrix $\Sigma$ whether or not there is a small amount of noise $R_{n}$ injected into the sum.
+The following lemma states that an analogous result is also true for the analogous (multiplier) bootstrap quantity under the relevant bootstrap law.
 {restatable}
@@
 $\sup_{t\in\mathbb{R}}\left|P^{B}\left(\hat{S}_{n,[k]}^{B}\leq t\right)-P\left(N(0,\Sigma)_{[k]}\leq t\right)\right|\xrightarrow[]{P}0.$
-Using these lemmas, we may now state two corollaries, which allow for studentization.
-First, define $\Lambda:=\text{diag}(\Sigma)$, as well as $\Sigma_{0}:=\Lambda^{-1/2}\Sigma\Lambda^{-1/2}$.
-Let $\widehat{\Sigma}_{jj}:=n^{-1}\sum_{i=1}^{n}\left(X_{ij}-\bar{X}_{n,j}\right)^{2}$ and let $\widehat{\Lambda}:=\operatorname{diag}\left\{\widehat{\Sigma}_{11},\ldots,\widehat{\Sigma}_{pp}\right\}$ be an estimator of the asymptotic variances $\Lambda$.
+Using these lemmas, we may now state two propositions that allow for studentization, recognizing that studentization with estimated variances can be cast as analyzing a particular $S_{n}$ under estimation error.
+We define the diagonal matrix of asymptotic variances $\Lambda:=\text{diag}(\Sigma)$, as well as the corresponding correlation matrix $\Sigma_{0}:=\Lambda^{-1/2}\Sigma\Lambda^{-1/2}$.
+Let plug-in estimates be given by $\widehat{\Sigma}_{jj}:=n^{-1}\sum_{i=1}^{n}\left(X_{ij}-\bar{X}_{n,j}\right)^{2}$ and $\widehat{\Lambda}:=\operatorname{diag}\left\{\widehat{\Sigma}_{11},\ldots,\widehat{\Sigma}_{pp}\right\}$.
 {restatable}
-[High-dimensional CLT for the small $k$-max studentized coordinate]corollaryCorhd If Assumptions M and R hold, and $B_{n}=O(1)$ and $k$ is fixed (i.e., does not grow with $n,p$), then, by application of Lemma 3.3, as $n,p\to\infty$
+[High-dimensional CLT for the small $k$-max studentized coordinate]propositionCorhd If Assumptions M and R hold, and $B_{n}=O(1)$ and $k$ is fixed (i.e., does not grow with $n,p$), then, by application of Lemma 3.3, as $n,p\to\infty$
 $\sup_{t\in\mathbb{R}}\left|P\left(\left(\hat{\Lambda}^{-1/2}S_{n}\right)_{[k]}\leq t\right)-P\left(N(0,\Sigma_{0})_{[k]}\leq t\right)\right|\to 0.$
+This proposition validates that the relevant estimation error $R_{n}$ in the studentization setting is indeed small enough in the sense required by Lemma 3.3 to lead to a standard high-dimensional CLT based on the relevant (asymptotic) correlation matrix.
+The next proposition proves the analogous result for the bootstrap case.
 {restatable}
-[High-dimensional bootstrap for the small $k$-max studentized coordinate]corollaryCorhdboot If Assumptions M and R hold, and $B_{n}=O(1)$ and $k$ is fixed (i.e., does not grow with $n,p$), then, by application of Lemma 3.3, as $n,p\to\infty$
+[High-dimensional bootstrap for the small $k$-max studentized coordinate]propositionCorhdboot If Assumptions M and R hold, and $B_{n}=O(1)$ and $k$ is fixed (i.e., does not grow with $n,p$), then, by application of Lemma 3.3, as $n,p\to\infty$
 $\sup_{t\in\mathbb{R}}\left|P^{B}\left(\left(\hat{\Lambda}^{-1/2}S_{n}^{B}\right)_{[k]}\leq t\right)-P\left(N(0,\Sigma_{0})_{[k]}\leq t\right)\right|\xrightarrow[]{P}0.$
-With these corollaries, it is immediate that the bootstrap quantiles of studentized statistics well approximate the studentized version of $S_{n}$.
-Also note that, in the setting where the data is not just i.n.i.d.
-but i.i.d., as is true for the use case of interest based on dictionary features $X_{i}$, $B_{n}=O(1)$ is assured.
+Combining these propositions, it is immediate that the bootstrap quantiles of studentized statistics well-approximate those of the studentized version of $S_{n}$.
+Further note that, in the setting where the data is not just i.n.i.d.
+but i.i.d., as is true for the use case of interest based on dictionary features extracted from an unstructured dataset obtained through i.i.d.
+sampling, $B_{n}=O(1)$ is assured.
+3.4 Generating and Evaluating Automatic Interpretations
+The preceding sections prove theoretical results necessary for developing $k$-FWER controlling statistical procedures that are valid in the high-dimensional setting encountered by this framework.
+From these procedures, the researcher obtains a selected set of rejected null hypotheses, or “discoveries.” Concretely, these discoveries are recorded as set of coordinates $\hat{J}\subseteq[p]$ indexing the rejected nulls.
+Though the machine learning theory that motivates sparse dictionary learning supposes that each hypothesis $\hat{\jmath}\in\hat{J}$ has single, human-interpretable meaning, the implementation of dictionary learning itself does not generate natural language descriptions of what each human-interpretable meaning may be.
+Human judgment alone could be deployed as a means of giving natural language descriptions to dictionary features, e.g., the researcher (or research assistants) could inspect many of the text examples that activate (highly) for a given dictionary feature in order to arrive at a plausible English language description for the concept signaled by that feature.
+However, such a process suffers from (1) problems related to motivated reasoning and data snooping (e.g., the researcher could simply come up with a feature description that is reverse engineered from a conclusion they wanted to find in the data); and (2) problems related to the scale of interpretation and associated interpretation effort/cost (e.g., there may be many hundreds of features discovered, all requiring interpretation; interpreting all $p$ features would require even more labor).
+A dictionary feature interpretation methodology that succumbs to neither of these problems is that of an “automatic interpretation” pipeline.
+The idea of an “autointerp” pipeline dates back to at least Bills et al.
+[2023], and seeks to automate the process of creating natural language descriptions for any set of dictionary features by leveraging a dedicated “explainer LLM” which can reason about feature interpretations cheaply at scale when provided with sufficient information.
+There is a large design space of possible autointerp pipelines, and many variants of autointerp methods have become popular in the machine learning interpretability literature.
+A particularly common autointerp generation strategy pioneered in Bricken et al.
+[2023] is similar to the human-driven procedure outlined above: pass several examples of texts that activate highly on a particular feature to an explainer LLM—possibly annotating these exemplar texts with information about where and how strongly the feature activates—along with a prompt that details the setting and asks for an English description of what the feature activation pattern in the presented texts indicates about the possible meaning of the feature.
+We adopt such an autointerp generation strategy for this framework, using text samples from the researcher’s unstructured dataset to make generated descriptions as relevant as possible to the unstructured data distribution of interest $P$.
+That is, generation is “localized” to the distribution $P$, unlike off-the-shelf dictionary features for pretrained models, which were learned on a different distribution of data (e.g., some large and opaque swath of the internet).
+The specifics of this generation strategy are detailed in Appendix Section 6.3.
+Example 1 (Continued).
+The null hypothesis of no average treatment effect is rejected for dictionary feature 4.
+Using the raw SAE feature activations (necessary to compute in the process of evaluating the Dict function), the researcher is able to rank text responses for experimental participants by largest activations for feature 4 on any token (“max pooling”).
+The researcher takes the top 10 largest activating text examples and annotates each of them with brackets delimiting the highest activating token in each response; these exemplar responses are then passed to an explainer LLM that reasons about and then outputs the concise English phrase “discussion of politics” to describe the concept feature 4 likely maps to (based on the activation patterns in the presented texts).
+Autointerp pipelines ideally not only specify how to generate descriptions for dictionary features, but also evaluate the quality of these descriptions.
+We now consider a new statistical formalization of a popular autointerp evaluation technique called “detection scoring.” The fundamental principle of detection scoring is that “an interpretation should serve as a binary classifier distinguishing activating from nonactivating contexts” [Paulo et al., 2024].
+In other words: description of a dictionary feature can be treated like a classification problem, one where the natural language description itself (incorporated in a LLM prompt) is the classifier, the input is a specific text example, and the label to be predicted is derived from the ground truth of the presence or absence of a feature’s activation in that text example (which is known to the researcher).
+Under this framing, binary classifier performance metrics like accuracy, precision, and recall characterize the quality of the description.
+To newly formalize this autointerp strategy, consider a “held-out” unstructured dataset of size $m$ drawn independently from the same distribution of interest $P$.
+This held-out dataset may be created, for example, by partitioning an original unstructured dataset drawn i.i.d.
+from $P$ into two disjoint datasets via indices chosen at random, independently of all the data.
+Let us denote the set of indices of unstructured data points in the main sample as $\mathcal{I}^{\textrm{estim}}$, and those in the held-out dataset as $\mathcal{I}^{\textrm{eval}}$.
+Using the data points with indices in $\mathcal{I}^{\textrm{estim}}$, we extract exemplar unstructured data points to be used to create an autointerp explanation of feature $j$, which we denote $\hat{\eta}_{j}$.
+For all $i\in\mathcal{I}^{\textrm{eval}}$, we define the random variable $\hat{Y}_{ij}\in\{0,1\}$ as
+$\hat{Y}_{ij}:=\texttt{CLS}(Z_{i},\hat{\eta}_{j}).$
+The function $\texttt{CLS}(\cdot)$ is a non-stochastic classification prompt fed both an unstructured data point $Z_{i}$ and the learned autointerp explanation $\hat{\eta}_{j}$.
+The $\hat{Y}_{ij}$ is the classifier output, a prediction of whether or not the unstructured data point $Z_{i}$ possesses a feature with concept description $\hat{\eta}_{j}$.
+[^12] Note that, even if a sample splitting implementation of this autointerp evaluation strategy is pursued, cross-fitting is inappropriate, because there is little reason to think, given the sample sizes available to researchers, that $\hat{\eta}_{j}$ would be stable across folds.
+As such, when sample splitting is employed, an uneven split is recommended, e.g., only 10% of the data is used for the held-out evaluation sample.
+For the purposes of evaluating autointerp descriptions, we define the “accuracy score” (or “A-score”) estimand as
+$\theta_{j}^{\textrm{acc}}(\hat{\eta}_{j};P):=E_{P}\left[S_{ij}\mid\hat{\eta}_{j}\right]=P(\{Y_{ij}=\hat{Y}_{ij}\}\mid\hat{\eta}_{j})=P\left(\{Y_{ij}=\texttt{CLS}(Z_{i},\hat{\eta}_{j})\}\mid\hat{\eta}_{j}\right)$
+where $S_{ij}:=\mathbf{1}\{Y_{ij}=\hat{Y}_{ij}\}$.
+Conditioning on $\hat{\eta}_{j}$ in the estimand of interest should be thought of as evaluating a fixed autointerp description.
+Importantly, as the notation emphasizes, the $\theta_{j}^{\textrm{acc}}(\hat{\eta}_{j};P)$ tells us exactly the quality of the description $\hat{\eta}_{j}$ for the population of interest $P$, unlike other autointerp evaluation methods for which the underlying distribution that implicitly defines the evaluation score is highly opaque (c.f., simulation scoring).
+We can easily construct a plug-in estimator of the autointerp accuracy (A-score) estimand; its unbiasedness, consistency, and asymptotic normality under the $\hat{\eta}_{j}$-conditional law are established in the following proposition.
+{restatable}
+[Conditional inference on A-score]propositionAscoreest Let $m:=|\mathcal{I}^{\textrm{eval}}|$. Define the A-score estimator for dictionary feature $j$ as
+$\hat{\theta}_{j}^{\textrm{acc}}:=\frac{1}{m}\sum_{i\in\mathcal{I}^{\textrm{eval}}}S_{ij}=\frac{1}{m}\sum_{i\in\mathcal{I}^{\textrm{eval}}}\mathbf{1}\{Y_{ij}=\hat{Y}_{ij}\}=\frac{1}{m}\sum_{i\in\mathcal{I}^{\textrm{eval}}}\mathbf{1}\{Y_{ij}=\operatorname{\texttt{CLS}}(Z_{i},\hat{\eta}_{j})\}.$
+Assume that $P(S_{ij}=1\mid\hat{\eta}_{j})\in[\delta,1-\delta]$ for some $\delta$ bounded away from 0 and 1, almost surely.
+Then we have that $E_{P}\left[\hat{\theta}_{j}^{\textrm{acc}}\mid\hat{\eta}_{j}\right]=\theta_{j}^{\textrm{acc}}(\hat{\eta}_{j};P)$ almost surely and as $m\to\infty$
+$\sup_{t\in\mathbb{R}}\left|P\left(\sqrt{m}(\hat{\theta}_{j}^{\textrm{acc}}-\theta_{j}^{\textrm{acc}}(\hat{\eta}_{j};P))\leq t\mid\hat{\eta}_{j}\right)-P(N(0,\operatorname{Var}(S_{ij}\mid\hat{\eta}_{j}))\leq t\mid\hat{\eta}_{j})\right|\xrightarrow[]{a.s.}0.$
+Example 1 (Continued).
+The null hypothesis of no average treatment effect is rejected for dictionary feature 4, and the autointerp generation strategy produced by the researcher describes feature 4 as activating on “discussion of politics.” To evaluate the quality of this description, the researcher runs a LLM-based classifier over a held-out dataset of experimental participant text responses; the LLM classifier is given the prompt “Does the following text example indicate discussion of politics?
+Return 1 or 0 for yes or no:” followed by a text response from the held-out dataset.
+The researcher then uses the output of the LLM classifier to compute the A-score, which measures the accuracy of this autointerp description-based classifier at predicting the presence of feature 4 in the held-out dataset.
+Towards better understanding the information communicated to the researcher by the A-score, it is useful to consider the function $\texttt{CLS}^{*}(\cdot)$, which we can define implicitly as
+$\bar{\theta}_{j}^{\textrm{acc}}(\hat{\eta}_{j};P):=\sup_{f\in\mathcal{F}}P\left(\{Y_{ij}=f(Z_{i},\hat{\eta}_{j})\}\mid\hat{\eta}_{j}\right)=P\left(\{Y_{ij}=\texttt{CLS}^{*}(Z_{i},\hat{\eta}_{j})\}\mid\hat{\eta}_{j}\right),$
+where $\mathcal{F}$ is the space of all viable classification prompts and LLMs able to execute them.
+That is, $\texttt{CLS}^{*}(\cdot)$ is the “oracle” detection scoring model setup, which when evaluated produces the highest interpretative accuracy score possible under a given description $\hat{\eta}_{j}$ under distribution $P$.
+Arguably, $\bar{\theta}_{j}^{\textrm{acc}}(\hat{\eta}_{j};P)$ is more informative to the researcher than $\theta_{j}^{\textrm{acc}}(\hat{\eta}_{j};P)$ about the quality of $\hat{\eta}_{j}$, as $\bar{\theta}_{j}^{\textrm{acc}}(\hat{\eta}_{j};P)$ purely quantifies the quality of $\hat{\eta}_{j}$, without the wedge introduced by classification failures stemming from $\texttt{CLS}(\cdot)$ itself.
+Because any feasible choice of $\texttt{CLS}(\cdot)$ yields
+$P\left(\{Y_{ij}=\texttt{CLS}(Z_{i},\hat{\eta}_{j})\}\mid\hat{\eta}_{j}\right)\leq P\left(\{Y_{ij}=\texttt{CLS}^{*}(Z_{i},\hat{\eta}_{j})\}\mid\hat{\eta}_{j}\right)$
+one can reasonably interpret $\hat{\theta}_{j}^{\textrm{acc}}$ as a conservative (under)estimate of the more informative quantity $\bar{\theta}_{j}^{\textrm{acc}}(\hat{\eta}_{j};P)$.
+We may also define estimands for autointerp precision and recall (the “P-score” and “R-score,” respectively):
+$\theta_{j}^{\textrm{prec}}(\hat{\eta}_{j})$ $:=P\left(Y_{ij}=1\mid\hat{Y}_{ij}=1,\hat{\eta}_{j}\right)=\frac{E[\mathbf{1}\{Y_{ij}=1\}\mathbf{1}\{\hat{Y}_{ij}=1\}\mid\hat{\eta}_{j}]}{E[\mathbf{1}\{\hat{Y}_{ij}=1\}\mid\hat{\eta}_{j}]},$
+$\theta_{j}^{\textrm{rec}}(\hat{\eta}_{j})$ $:=P\left(\hat{Y}_{ij}=1\mid Y_{ij}=1,\hat{\eta}_{j}\right)=\frac{E[\mathbf{1}\{Y_{ij}=1\}\mathbf{1}\{\hat{Y}_{ij}=1\}\mid\hat{\eta}_{j}]}{E[\mathbf{1}\{Y_{ij}=1\}\mid\hat{\eta}_{j}]}.$
+As above, these estimands may be consistently estimated with appropriate asymptotically valid confidence intervals under the $\hat{\eta}_{j}$ conditional law using the natural plug-in estimators and the delta method.
+A similar “lower bound” interpretation applies to these estimators.
+If a researcher is interested in inference on the best performing autointerp description across multiple generation prompts or strategies—say among a set $\{\hat{\eta}_{j,1},\hat{\eta}_{j,2},\dots\}$ for any $j$ of interest—then the evaluation performance should be estimated with appropriate “inference on winners” corrections such as those discussed in Andrews et al.
+[2024].
+Further, recall that we are not typically interested in autointerp evaluation for a fixed $j$, but instead for a particular data-driven choice $\hat{\jmath}\in\hat{J}$ suggested by a high-dimensional selective inference procedure.
+[^13] Autointerpretation typically incurs non-negligible compute costs, and, as such, researchers will often only be interested in forming meaningful autointerpretations for the subset of discovered concepts.
+However, so long as discovery is also performed using the main estimation sample indexed by $\mathcal{I}^{\textrm{estim}}$, the researcher may treat each $\hat{\jmath}$ as fixed for the purposes of evaluation inference.
+Formally, for all $j\in[p]$,
+$E_{P}\left[S_{i\hat{\jmath}}\mid\hat{\eta}_{\hat{\jmath}},\hat{\jmath}=j\right]=E_{P}\left[S_{ij}\mid\hat{\eta}_{j},\hat{\jmath}=j\right]=E_{P}\left[S_{ij}\mid\hat{\eta}_{j}\right]$
+where the last equality follows because, under the proposed independently held-out data (sample splitting) scheme,
+$\mathbf{1}\{Y_{ij}=\texttt{CLS}(Z_{i},\hat{\eta}_{j})\}\perp\!\!\!\perp\hat{\jmath}]\mid\hat{\eta}_{j}.$
+A well-defined and intuitive quantitative method for evaluating the quality of feature descriptions is important in the context of the proposed framework, as it is these descriptions on which the scientific conclusions of discovery ultimately rest.
+By formulating autointerp evaluation as a statistical inference problem, the importance of evaluation with respect to the population of interest $P$ becomes salient (as opposed to some other distribution implicitly defined through some other collection of unstructured data, common in most other autointerp evaluation methods); the need for sample splitting becomes clear; the role for confidence intervals becomes legible by appeal to sampling uncertainty (c.f., Miller [2024]); and we further unlock useful connections to the literature on post-selection inference, such that we may discipline settings evaluating a multiplicity of interpretations (c.f., Andrews et al.
+[2024]).
+It is worth noting that this autointerp evaluation strategy can also serve as a general-purpose description evaluation technology, e.g., for carefully crafted human descriptions of feature activations in settings where there is a small number of discoveries.
+The scientific benefit of evaluation in such a setting is also clear: even if there is concern that human-made descriptions are reverse-engineered to support a presupposed conclusion, any description (human or machine made, motivated or not) that is not supported by the data will be surfaced through the evaluation scores.
+In this way, description evaluation can be seen as even more critical when automatic methods are not employed by the researcher to generate descriptions.
+3.5 Pseudocode Algorithm
+We now state the entire framework for interpretable discovery from unstructured data as a single pseudocode algorithm.
+For simplicity, we focus on an implementation with text data, sample splitting, two-sided testing, one-step inference, and autointerp evaluation via accuracy scoring (A-scores).
+Algorithm 1 Interpretable Discovery from Text Data with $k$-FWER Control (Simplified)
+1:Data $\{(W_{i},Z_{i})\}_{i=1}^{n+m}$ drawn i.i.d.
+from distribution $P$; dictionary function $\mathtt{Dict}:\mathcal{Z}\to\{0,1\}^{\tilde{p}}$; map $h:\mathcal{W}\times\{0,1\}\to\mathbb{R}$; level $\alpha$; error control $k$; explainer LLM $\mathcal{E}$; eval dataset size $m$; bootstrap iterations $\beta$.
+[…]
+3:Split sample
+4:Randomly split $[n+m]$ into $\mathcal{I}^{\mathrm{estim}}$ and $\mathcal{I}^{\mathrm{eval}}$ with $|\mathcal{I}^{\mathrm{estim}}|=n$ and $|\mathcal{I}^{\mathrm{eval}}|=m$.
+5:Create concept dictionaries
+6:for $i=1,\dots,n+m$ do
+7:  $Y_{i}\leftarrow\mathtt{Dict}(Z_{i})\in\{0,1\}^{\tilde{p}}$.
+8:end for
+9:Drop degenerate features
+10:Initialize $\mathcal{J}\leftarrow\emptyset$.
+11:for $j=1,\dots,\tilde{p}$ do
+12:  If $\exists\,i\text{ such that }Y_{ij}\neq 0$, $\mathcal{J}\leftarrow\mathcal{J}\cup\{j\}$.
+13:end for
+[…]
+15:Create feature-level estimates and t-stats
+16:for $j\in\mathcal{J}$ do
+17:  For $i\in\mathcal{I}^{\mathrm{estim}}:X_{ij}\leftarrow h(W_{i},Y_{ij})$.
+18:  $\hat{\theta}_{j}\leftarrow\frac{1}{n}\sum_{i\in\mathcal{I}^{\mathrm{estim}}}X_{ij};\hat{\Sigma}_{jj}\leftarrow\frac{1}{n}\sum_{i\in\mathcal{I}^{\mathrm{estim}}}(X_{ij}-\hat{\theta}_{j})^{2}$.
+19:end for
+[…]
+21:High-dimensional selective inference
+22:Critical value $\hat{c}_{n}(1-\alpha,k)\leftarrow 1-\alpha\text{ quantile of }|S_{n}^{B}|_{[k]}$ under bootstrap law, approx.
+using $\beta$ multiplier bootstrap iterations with $\{X_{i}\}_{i\in\mathcal{I}^{\textrm{estim}}}$.
+[…]
+24:For $j\in[p]$: gen.
+sim.
+CI $\mathcal{C}_{j}\leftarrow\left[\hat{\theta}_{j}-\hat{c}_{n}(1-\alpha,k)\sqrt{\frac{\hat{\Sigma}_{jj}}{n}},\hat{\theta}_{j}+\hat{c}_{n}(1-\alpha,k)\sqrt{\frac{\hat{\Sigma}_{jj}}{n}}\right]\in\mathbb{R}^{2}$.
+25:Autointerpretation
+26:Intialize CLS using $\mathcal{E}$ and fixed classification prompt.
+27:for $\hat{\jmath}\in\hat{J}$ do
+28:  Generate descriptions $\hat{\eta}_{\hat{\jmath}}\leftarrow\mathcal{E}(\{Z_{i}\}_{i\in\mathcal{I}^{\textrm{estim}}})$ (e.g., implemented per App. Sec. 6.3).
+29:  for $i\in\mathcal{I}^{\mathrm{eval}}$ do
+30:   $\hat{Y}_{i\hat{\jmath}}\leftarrow\mathtt{CLS}(Z_{i},\hat{\eta}_{\hat{\jmath}})\in\{0,1\};S_{i\hat{\jmath}}\leftarrow\mathbf{1}\{Y_{i\hat{\jmath}}=\hat{Y}_{i\hat{\jmath}}\}$.
+31:  end for
+32:  A-score $\hat{\theta}_{\hat{\jmath}}^{\mathrm{acc}}\leftarrow\frac{1}{m}\sum_{i\in\mathcal{I}^{\mathrm{eval}}}S_{i\hat{\jmath}}$.
+[…]
+34:end for
+35:Output: $\{(j,\hat{\theta}_{j},\mathcal{C}_{j}):j\in[p]\}$ and $\{(\hat{\jmath},\hat{\eta}_{\hat{\jmath}},\hat{\theta}_{\hat{\jmath}}^{\mathrm{acc}}):\hat{\jmath}\in\hat{J}\}$.
 4 Empirical Applications
@@
 4.1 Bursztyn et al. [2023]
-Bursztyn et al.
+Mirroring Example 1, Bursztyn et al.
 [2023] study how the provision of “social cover” affects willingness to publicly dissent to socially stigmatized causes, and the perception of this dissent.
@@
 [2023] run an information treatment experiment (Experiment 2) in which participants are told they have been matched with another (fictional) respondent that chose to join a campaign to defund the police (a plausible expression of dissent in liberal American politics at the time the experiment was conducted), and then show the participant a tweet that the matched respondent is said to have agreed to publicly post.
-This tweet has been randomized to either include social cover or no social cover, namely whether or not the tweet indicated that the matched respondent joined the campaign after reading an evidence-based article in support of it (the no social cover condition) or before (the social cover condition).
+This tweet has been randomized to either include social cover or no social cover, namely whether or not the tweet indicated that the matched respondent joined the campaign before reading an evidence-based article in support of it (the no social cover condition) or after (the social cover condition).
 [^14] See Online Appendix Figure B.3 of Bursztyn et al.
@@
 An important outcome collected by this experiment is the participant’s open-ended text response to the question “Why do you think your matched respondent chose to join the campaign to oppose defunding the police?” This open-ended text response is meant to capture the causal effect of social cover on the perception of dissent.
-As such, this is a setting where discovery is of interest; ideally, to form a holistic understanding of how social cover affects perception of dissent, we do not want to pre-specify what aspects of perception we are interested in, and want to discover any interpretable systematic differences that exist in the text responses across treatment and control groups.
+As such, this is a setting where discovery is of interest: ideally, to form a holistic understanding of how social cover affects perception of dissent, we do not want to pre-specify what aspects of perception we are interested in, and want to discover any interpretable systematic differences that exist in the text responses across treatment and control groups.
 In order to make progress on this analysis of unstructured data, Bursztyn et al.
@@
 [2023] contains the top ten characteristic phrases in each condition based on the $\chi^{2}$ index.
-Can we make more (and more interpretable) discoveries with the framework of this paper?
-To investigate, we implement a Dict function based on the Gemma Scope autoencoders for Google’s Gemma 2 2B [Lieberum et al., 2024], and specifically the SAE trained on the residual stream of layer 12.777It is recommended to work with SAEs trained on the middle layer of a LLM to capture coarser concepts and features, which is plausibly more suitable for analyzing broad themes in text.
-We pull the Gemma Scope project’s autointerpretation descriptions of these features from Neuronpedia, [^8] See https://www.neuronpedia.org/gemma-scope for more.
-though we also implement a custom, “local” autointerpretation pipeline based on the best practices of Paulo et al.
-[2024], discussed further in Appendix Section 6.2.
-These are mainstream choices for a SAE and LLM pairing in machine learning interpretability research.
-This Dict function yields a dataset $\{X_{ij}\}_{i\in[n],j\in[p]}$, where $n=1033$ and p=12,005p=12,\!005 (non-degenerate features). The hypotheses of interest are, for all $j\in[p]$,
-$H_{0,j}:E[Y_{ij}(1)-Y_{ij}(0)]=E[H_{i}Y_{ij}]:=E[X_{ij}]=0,\quad H_{i}=\frac{D_{i}-\pi}{\pi(1-\pi)}.$
-for $\pi=0.5$, and where $\{D_{i}=1\}$ indicates the social cover condition.
-As a baseline, we will implement a simple one-step version of the statistical procedure in Theorem R, which will attain a lower bound on the power achieved by the step-wise procedure.
+Can we make more (and more interpretable) discoveries with the framework proposed in this paper?
+To investigate, we implement a Dict function based on the Gemma Scope autoencoders for Google’s Gemma 2 2B LLM [Lieberum et al., 2024], and specifically the SAE trained on the residual stream of layer 12.171717It is a well-documented observation that SAEs trained on the middle layer of a LLM seem to capture “coarser” concepts and features (see dicussion in, e.g., Templeton et al.
+[2024]), which is plausibly more suitable for analyzing broad themes in text.
+We implement an autointerpretation pipeline as discussed in Section 3.4 using as an explainer LLM the lightweight, open-source reasoning model GPT-OSS 20B [OpenAI, 2025].
+A held-out evaluation dataset is created by sample splitting, using 10% of the original dataset in Bursztyn et al.
+[2023].
+The hypotheses of interest are, for all $j\in[p]$,
+$H_{0,j}:E[Y_{ij}(1)-Y_{ij}(0)]=E\left[\frac{W_{i}-\pi}{\pi(1-\pi)}Y_{ij}\right]:=E[X_{ij}]=0,$
+for $\pi=0.5$, and where the event $\{W_{i}=1\}$ indicates the social cover condition.
+This yields a dataset $\{X_{ij}\}_{i\in[n],j\in[p]}$, where $n=930$ and $p=11901$ (non-degenerate features).
+We implement the simple one-step version of the statistical procedure in Theorem R as per Algorithm 1, which attains a lower bound on the power achieved by the step-wise procedure.
 We may first consider whether or not, as anticipated, simply controlling the FWER would be too conservative for the purposes of discovery in this setting.
-To explore this, we set $k=1$ and $\alpha=0.05$ in the procedure based on Theorem R (equivalently, we implement the high-dimensional FWER controlling procedure of Belloni et al.
-[2018]), and we obtain only two significant discoveries: a 13.2 percentage point causal effect on the presence of dictionary feature 3518 (DF-3518) and a 13.4 percentage point causal effect on the presence of dictionary feature 3426 (DF-3426).
-The Gemma Scope autointerprations of these features paint a picture that corroborates the original findings of Bursztyn et al.
-[2023]: DF-3426 is described as activating on “references to articles” and DF-3426 on “references to news articles and reports.” However, we only discover two features out of over 12,00012,\!000, and not much new insight relative to the original analysis—even if the interpretation of the features discovered was transparent and automatic, by construction.
-In Table 1, we make discoveries based on the procedure of Theorem R now setting $k=5$, still with $\alpha=0.05$, and this time using local autointerpretation descriptions per the method Appendix Section 6.2.
-Table 1: Discoveries for Experiment 2 of Bursztyn et al. (2023), 5-FWER Control at $\alpha=5\%$
-Feature $\widehat{\text{ATE}}$ Description (“Feature activates highly on…”) t-stat
-306 0.099 source that claims police defunding increases crime 4.103
-1392 0.118 partial or incomplete reading of an article 3.910
-1622 0.072 mentions of a prestigious author or institution (e.g., “Princeton professor”) 4.045
-1992 0.095 mentions of an article 4.286
-3298 0.126 article about defunding the police and its impact on violent crime 4.223
-3426 0.134 mentions of an “article” or article reference 4.803
-3518 0.132 articles linking police defunding to increased violent crime 5.020
-4320 0.081 article or report content 3.999
-8486 0.085 defunding‑the‑police political discourse 3.918
-8902 0.079 reading an article that influences a person’s beliefs or viewpoints 4.198
-12287 -0.130 modal verbs expressing uncertainty or possibility 3.890
-Note: Descriptions generated using the method described in Appendix Section 6.2, using GPT-OSS 20B [OpenAI, 2025]. For bootstrapping, 10,00010,\!000 iterations are performed.
-As can be seen, even with control of $k$-FWER at $k=5$, we obtain 11 rejections of the null, or 11 discoveries.
-By guarantee of Theorem R, the probability that 5 or more of these 11 discoveries is false is less than 5%.
+To explore this, we set $k=1$ and $\alpha=0.05$ (equivalently, we implement the high-dimensional FWER controlling procedure of Belloni et al.
+[2018]), and we obtain only two significant discoveries: a 13.2 percentage point causal effect on the presence of dictionary feature 3518 and a 13.4 percentage point causal effect on the presence of dictionary feature 3426.
+The autointerprations of these features paint a picture that corroborates the original findings of Bursztyn et al.
+[2023]: feature 3426 is described as activating on mentions of “defunding the police associated with increased violent crime” and feature 3518 on “mention of an article as a source of evidence.” However, we only discover two features out of approximately $12000$, and not much new insight relative to the original analysis—even if the discovery and interpretation of these features was transparent, inexpensive, and automatic.
+[^18] The off-the-shelf Gemma Scope autointerpretations echo the same: feature 3426 is described as activating on “references to news articles and reports” and feature 3518 on “references to articles.”
+Figure 1: Discoveries for Experiment 2 of Bursztyn et al. (2023)
+In Figure 1, we make discoveries now setting $k=5$, still with $\alpha=0.05$.
+As can be seen, even with control of $k$-FWER at small $k=5$, we obtain 9 rejections of the null, or 9 discoveries.
+By (large sample) guarantee of Theorem R, the probability that 5 or more of these 9 discoveries is false is less than 5%.
 The previous features discovered with the FWER controlling procedure appear, as well as others that represent concepts both absent from the FWER controlled analysis and the original analysis in Bursztyn et al.
 [2023].
-Though there are features that appear to activate on redundant concepts, in totality these discoveries contribute to entirely new characterizations of the causal effect of social cover on dissent.
-For example, DF-1622 responds to “mentions of a prestigious author or institution” and increases by 7 percentage points in the treatment group, indicating that, as Bursztyn et al.
-[2023] posit, the social cover mechanism plausibly relies on the perceived credibility of the article cited.
-DF-1392 increases by 12 percentage points in the treatment group, and responds to discussion of “partial or incomplete reading of an article,” suggesting that that social cover in this experiment may have a negatively valenced effect related to perceptions of ignorance.
-DF-12287 responds to “modal verbs expressing uncertainty or possibility,” and its probability of activation is decreased by 13 percentage points in the treatment group.
-This supports the idea that the causal effect of not having social cover is, to first order, generating uncertainty and speculation about dissent, as opposed to an alternative where participants coalesce around a single narrative in the absence of cover.
+Though there are features that appear to activate on redundant concepts, in totality these discoveries contribute to a few entirely new characterizations of the causal effect of social cover on dissent.
+For example, feature 1622 responds to “Prestigious author source reference” and increases by nearly 8 percentage points in the treatment group, indicating that—exactly as Bursztyn et al.
+[2023] posit—the social cover mechanism plausibly relies on the perceived credibility of the article cited.
+Feature 1355 responds to “modal verbs expressing uncertainty or possibility,” and its probability of activation is decreased by nearly 12 percentage points in the treatment group.
+This supports the idea that the causal effect of not having social cover is, to first order, generating uncertainty and speculation about dissent, as opposed to an alternative scenario where participants coalesce around a single narrative in the absence of cover.
 This insight is also congruous with the observation that few large negatively signed causal effects are discovered.
-Importantly, all of this analysis was performed automatically, in a handful of minutes on a single A100 GPU in Google Colab.
+The A-scores for all discoveries range from 0.98 to 0.66, indicating that some descriptions of discoveries deserve more skepticism than others, and that there is much scope for improving autointerp generation beyond small, open-source reasoning LLMs with the prompting strategy of Appendix Section 6.3.
+Importantly, all of this analysis was performed automatically, in approximately one hour, on a single A100 GPU in Google Colab.
 Though this analysis was not preregistered, it is possible to cheaply and automatically replicate these results, and assess their sensitivity to different autointerpretation strategies, dictionary learning methods, and LLMs.
-Notably, the space of possible concepts was not pruned in any way prior to this analysis, nor were they weighted in some way towards topics related to articles; the analysis was automatic and without human intervention, with no room for motivated data mining.
+Notably, the space of possible concepts was not pruned in any way prior to this analysis, nor were they weighted in some way towards topics related to articles; the analysis was automatic and without human intervention, with no room for motivated data snooping.
 4.2 Stantcheva [2024]
-Stantcheva [2024] conducts surveys on representative samples of the United States population investigating attitudes towards inflation.
+Mirroring Example 2, Stantcheva [2024] conducts surveys on representative samples of the United States population investigating attitudes towards inflation.
 These surveys include open-ended text responses to a variety of questions, with the goal of discovering attitudes and opinions that the researcher should not (for concern about priming respondents) or could not (for lack of imagination) pre-specify.
@@
 [2024] for more discussion of the benefits of open-ended survey questions for understanding economic behavior.
-An important open-ended prompt that Stantcheva [2024] solicit an answer to is “High inflation is caused by…” In the original analysis, Stantcheva [2024] code open-ended responses with a keyword-based method described in Ferrario and Stantcheva [2022], in which the researcher creates a list of topics and associated keywords in a discretionary way, ranging from “manual to semi-supervised or unsupervised” means.
+An important open-ended prompt that Stantcheva [2024] solicit an answer to is “High inflation is caused by…” In the original analysis, Stantcheva [2024] codes open-ended responses with a keyword-based method introduced in Ferrario and Stantcheva [2022] in which the researcher creates a list of topics and associated keywords in a discretionary way, which may range from “manual to semi-supervised or unsupervised” means.
 The result of applying this analysis to the open-ended responses to the above prompt yields Figure 3 in Stantcheva [2024], which finds, among other things, that mentions of “Biden and the administration,” “Greed,” “Monetary policy,” “Fiscal policy,” “War and foreign policy,” “Demand vs supply,” and “Supply-side mechanisms (other than input prices)” appear in more than 5% of all responses.
@@
 Do we make more and new interpretable discoveries?
-To find out, we use the same models as in Section 4.1, but reduce the space of features by half, filtering out all features that, in the corpus on which the Gemma Scope SAEs were trained, had greater than median empirical token activation frequency.
+To find out, we use the same Dict function and explainer LLM as in Section 4.1, but reduce the space of features by half, filtering out all features that, in the corpus on which the Gemma Scope SAEs were trained, had greater than median empirical token activation frequency.
 This choice of dimensionality reduction is meant to screen out features that activate in many texts across domains, e.g., features related to grammatical aspects of text, which was handled implicitly in the previous causal analysis by virtue of differencing.
-(Ideally, such a dimensionality reduction choice would be preregistered, to prevent cherry-picking by filtering.) This yields a dataset $\{X_{ij}\}$ with $n=503$ and $p=3915$ non-degenerate features—still a very high-dimensional inference problem.
-We test the $p=3915$ hypotheses $H_{0,j}:E[X_{ij}]=E[Y_{ij}]=0$, i.e., we test the probability of a given feature activating in the population of open-ended responses to the above prompt in the United States.
-Specifically, by again using 5-FWER control with $\alpha=0.05$ based on the procedure of Theorem R, we make 818 interpretable discoveries.
-In order to better present these discoveries, using test inversion we produce simultaneous “generalized” confidence sets [Romano and Wolf, 2007], which asymptotically guarantee that the probability that $k$ or more estimands do not lie in the set is less than $\alpha$.
-In Table 2, we report the simultaneous confidence intervals for all discoveries with a lower confidence bound above 20%.
-Notably, we discover, automatically, with no manual pruning or interpretation required, many sensible commonplace themes related to the topic at hand: DF-9804 responds to economic terms, DF-14304 responds to economic terminology, and DF-13447 responds to mentions of inflation.
-However, we also recover many of the topics from the original analysis in Stantcheva [2024] without any human discretion or intervention: DF-4192 responds to discussion of monetary policy, DF-5719 responds to political commentary, and DF-14747 responds to discussion of supply-side issues.
-Undoubtedly, many other topics are recovered as we look deeper down the list of discoveries, beyond the top 11 out of 818.
-However, even in this top 11, we learn that there is a great deal of pure uncertainty expressed in these open-ended responses based on the activation of DF-4794; that macroeconomic indicators get frequently discussed by virtue of DF-11036; and that money supply and government spending and budgeting are often mentioned, per DF-104 and DF-8316, respectively.
-Table 2: Largest Discoveries for “High inflation is caused by…” in Stantcheva (2024)
-Feature Description (“Feature activates highly on…”) Sim. CI, Lower Sim. CI, Upper
-104 excess money supply 0.208 0.352
-4192 monetary policy and central‑bank terminology 0.217 0.363
-4794 expressions of uncertainty or hedging 0.210 0.355
-5719 political commentary about government or the economy 0.843 0.942
-8316 government budget or spending 0.279 0.433
-9804 economic terms related to the economy 0.268 0.420
-11036 macroeconomic indicators 0.212 0.357
-13447 inflation or rising prices 0.393 0.553
-13574 negative economic conditions 0.230 0.378
-14304 economic terminology 0.300 0.456
-14747 supply‑related terms 0.234 0.382
-Note: Generalized confidence intervals based on $k$-FWER control at $k=5$, $\alpha=5\%$.
-Descriptions generated using the method described in Appendix Section 6.2, using GPT-OSS 20B [OpenAI, 2025].
-For bootstrapping, 10,00010,\!000 iterations are performed.
-Once again, this procedure was cheap and fast to implement, and could be replicated and analyzed for sensitivity by any other researcher, quickly.
-Beyond filtering out greater than median activating dictionary features, no other choices were made to delimit the space of possible discoveries, or focus them in some way on the space of economic- or inflation-relevant concepts; the exact same models that yielded the results of Section 4.1 yielded the results in this section.
+(Ideally, such a dimensionality reduction choice would be preregistered, to prevent cherry-picking by filtering.) Again we split the sample randomly to form a held-out evaluation dataset consisting of 10% of the original dataset analyzed in Stantcheva [2024].
+We are interested in testing hypotheses $H_{0,j}:E[X_{ij}]:=E[Y_{ij}]=0$, i.e., we test the probability of a given feature activating in the population of open-ended responses to the above prompt in the United States.
+This yields a dataset $\{X_{ij}\}_{i\in[n],j\in[p]}$ with $n=453$ and $p=3826$ non-degenerate features—still a very high-dimensional inference problem.
+We now apply 5-FWER control with $\alpha=0.05$ based on the result in Theorem R per Algorithm 1, and we make 744 interpretable discoveries.
+In Figure 2, we report point estimates and simultaneous confidence intervals for all discoveries with a generalized simultaneous lower confidence bound above 20%.
+Notably, we discover, automatically, with no manual pruning or interpretation required, many sensible commonplace themes related to the topic at hand: feature 9804 responds to discussion of the economy; feature 13447 responds to mentions of inflation; and feature 13574 activates on broad discussion of negative economic conditions.
+However, we also recover many of the topics from the original analysis in Stantcheva [2024] without any human discretion or intervention: feature 4192 responds to discussion of monetary policy; feature 14747 responds to discussion of supply-side issues; feature 104 activates on discussion of excess and overabundance; feature 8316 responds to fiscal policy mentions; and feature 5719 responds to economic criticisms leveled at the current government.
+Undoubtedly, many other topics are recovered as we look deeper down the list of discoveries, beyond these 11 largest.
+However, even in this top 11, we learn that there is a great deal of pure uncertainty expressed in these open-ended responses based on the activation of feature 4794, and also that macroeconomic indicators get frequently discussed by virtue of feature 11036.
+Figure 2: Ranked Discoveries for “High inflation is caused by…” in Stantcheva (2024)
+Once again, this procedure was cheap and fast to implement, and could be quickly replicated and analyzed for sensitivity by any other researcher.
+Beyond filtering out greater than median activating dictionary features, no other choices were made to delimit the space of possible discoveries, or focus them in some way on the space of economic- or inflation-relevant concepts; the exact same explainer LLM and SAE that yielded the results of Section 4.1 yielded the results in this section.
 5 Conclusion
@@
 The framework proposed in this paper shows how new statistical procedures for high-dimensional multiple hypothesis testing, when combined with the latest innovations in interpretability methods for machine learning models, can facilitate open-ended, interpretable discovery at scale with higher practicality and higher fidelity than previously possible.
-Importantly, the automaticity of the proposed framework makes it subject to very few researcher degrees of freedom, making it resilient to cherry-picking and motivated data-mining without compromising the purpose of discovery in the first place.
-This is especially true if one couples this framework with minimal preregistration efforts, e.g., simply publicly declaring what dictionary learning model, LLM, and automatic interpretation method are to be used for the analysis of a text, and if any of the features are to be filtered on principled grounds.
-A reasonable possible criticism of the proposed framework is that the interpretations associated with dictionary learning methods are not credible.
-On this topic exactly, there has been an immense amount of attention from AI/ML research communities, which has, to date, primarily concluded that such dictionary features are useful and reliable for discovery, even if they do not uncover “atomic” or fundamental conceptual units in deep neural networks like LLMs [Leask et al., 2025, Peng et al., 2025].
-The best practices and methods for moving from features to feature descriptions is an open and important question in the literature, and choices made on this front certainly affect interpretations of discoveries made by the proposed framework.
-However, as the empirical examples of the previous section make clear, straightforward automatic interpretation methods—which are completely data-driven, replicable, and not subject to motivated human reasoning—yield sensible outputs, scale well with the size of selected sets, can be further scrutinized by humans to assess their validity, and can be perturbed or modified cheaply to assess sensitivity to autointerpretation prompts.
-Future research related to this framework would seek to apply quantitative methods for validating feature descriptions at scale as well [Paulo et al., 2024, Movva et al., 2025].
+It also illustrates how treating (automatic) feature interpretation evaluation as a formal statistical inference problem can yield several scientific benefits in the process of making discoveries.
+Importantly, the automaticity of the proposed framework makes it subject to very few researcher degrees of freedom, making it resilient to cherry-picking and motivated data snooping without compromising the purpose of discovery in the first place.
+This is especially true if one couples this framework with minimal preregistration efforts, e.g., simply publicly declaring what dictionary learning model, LLM, and automatic interpretation pipeline (explainer LLM and prompts) are to be used for the analysis, and if any of the features are to be filtered on principled grounds.
 The proposed framework is most naturally viewed as one tool of many in the empirical researcher’s toolkit for making discoveries from unstructured data.